### v2

# and hierarchical hyperbolicity

Abstract.
We call a central extension bounded if its Euler class is represented by a bounded cocycle. We prove that a bounded central extension of a hierarchically hyperbolic group (HHG) is still a HHG; conversely if a central extension is a HHG, then the extension is bounded, and under a further mild assumption the quotient is commensurable to a HHG. Motivated by questions on hierarchical hyperbolicity of quotients of mapping class groups, we therefore consider the general problem of determining when a quotient of a bounded central extension is still bounded, which we prove to be equivalent to an extendability problem for quasihomomorphisms. Finally, we show that quotients of the 4-strands braid group by suitable powers of a pseudo-Anosov are HHG, and in fact bounded central extensions of some HHG. We also speculate on how to extend the previous result to all mapping class groups.
1. Introduction
Central extensions $1\to K\to E\to G\to 1$ (which we often abbreviate with just $E$) are classified by their associated Euler class $\alpha_{E}\in H^{2}(G;K)$. We call a central extension bounded if its Euler class is bounded, i.e. it is represented by a bounded cocycle. One reason of interest is the fact that bounded central extensions are quasi-isometrically trivial [Gersten_B_is_QIT]. Another reason of interest comes from the study of hierarchically hyperbolic groups, see below.
In this paper we always consider central extensions with finitely generated kernel. We are interested in the following natural problem:
Problem 1.1.
Given a bounded central extension $E$, which of its quotient central extensions $\bar{E}$ are bounded?
By a quotient central extension here we mean that there exists a diagram as follows:
$1{1}K{K}E¯{\bar{E}}G¯{\bar{G}}1.{1.}\scriptstyle{\cong}$
Equivalently, $\bar{E}=E/N$, where $N$ is a normal subgroup of $E$ intersecting $K$ trivially.
We note in the other direction that if a central extension has a bounded quotient central extension, then it is bounded (Lemma 2.12). However, not all quotient extensions of a bounded central extension are themselves bounded. For instance, the Heisenberg group is a quotient central extension of $F_{2}\times\mathbb{Z}$ which is not bounded (see Example 2.13); in fact, any central extension can be realised as a quotient of a trivial extension of a free group (see Example 2.14).
By Proposition 2.9 (which is probably known to experts, and was recently obtained independently by Tao and Wan [tao2025properactionsfiniteproducts, Lemma 5.6]), boundedness of a central extension $E$ is equivalent to the existence of a quasihomomorphism $E\to K$ that is the identity on $K$. Quasihomomorphisms, in the sense of e.g. [fujikapo], are also relevant for quotient central extensions, and indeed we show that boundedness of a quotient central extension is equivalent to an extendability problem for a quasihomomorphism. An informal statement is as follows:
Proposition 1.2 (see Proposition 2.16).
Consider a bounded central extension
[…]
with finitely generated kernel, and a quotient central extension $\bar{E}=E/N$. Then $\bar{E}$ is bounded if and only if a certain quasihomomorphism on $\pi(N)$ extends to $G$.
This has further connection to bounded cohomology in degree 3, which we discuss in Section 2.
The second and third authors encountered Problem 1.1 while studying quotients of mapping class groups, in particular relating to their conjecture on hierarchical hyperbolicity of quotients of mapping class groups [MS_rigidity_MCG, Question 3]. Hierarchically hyperbolic groups (HHG), as first defined by Behrstock, Hagen, and Sisto in [HHS_I], provide a common framework for, among others, mapping class groups of surfaces, most cubulated groups, fundamental groups of three-manifolds, and extra-large Artin groups [hagensusse, HRSS_3manifold, ELTAG], and are therefore amenable to tools from both low-dimensional topology and the world of CAT(0) cube complexes. Showing that a given group is hierarchically hyperbolic yields a lot of information about it ([quasiflat, HHP:coarse, ANS:UEG, HHL:proximal, DMS:stable, FarrellJones] is a highly non-exhaustive list). Hence it is natural to explore which group-theoretic procedures the class of HHG is closed under. These include taking graph products [graphprod], relative hyperbolicity [HHS_II], many graphs of groups [berlairobbio], and several quotients [BHMS, short_HHG:II, random_quot]. In this direction, we characterise which central extensions preserve hierarchical hyperbolicity:
Theorem 1.3 (see Theorem 3.11).
Let $1\to K\to E\to G\to 1$ be a central extension with finitely generated kernel, and suppose that $G$ is a hierarchically hyperbolic group. Then $E$ is a hierarchically hyperbolic group if and only if the extension is bounded.
We postpone to Section 4 the details of how Theorem 1.3 relates to hierarchical hyperbolicity of quotients of mapping class groups. In the other direction, under a further mild assumption, we prove that a central quotient of a HHG is commensurable to a HHG:
Theorem 1.4 (see Theorem 3.14).
Let $1\to K\to E\to G\to 1$ be a central extension with finitely generated kernel. If $E$ is a hierarchically hyperbolic group, then the extension is bounded.
If moreover $E$ has a HHG structure with clean containers, there exists a finite-index subgroup $E^{\prime}\leq E$ containing $K$ such that $E^{\prime}/K$ is a hierarchically hyperbolic group.
The hypothesis that $E$ has clean containers is mild and natural, as explained in Definition 3.5.
Since $G$ is a finite-index overgroup of a HHG, it inherits the structure of a hierarchically hyperbolic space, given by the quasi-isometric inclusion $E^{\prime}/K\hookrightarrow G$. However, even assuming that $E$ has clean containers, one cannot hope that $G$ itself is a hierarchically hyperbolic group: in Remark 3.15 we argue that the trivial central extension $E=\mathbb{Z}\times G$ of the $(3,3,3)$-triangle group $G$ is HHG, while $G$ is not by [PS_unbounded].
Finally, inspired by the setup explained in Section 4, we show that many quotients of the braid group $B_{4}$ on 4 strands are bounded central extensions. Recall here that $B_{4}$ is a central extension of a finite-index subgroup of the mapping class group of the five-holed sphere. The quotients we consider here are quotients by powers of pseudo-Anosovs, as studied for finite-type surfaces for instance in [DGO, hhs_asdim].
Theorem 1.5 (see Theorem 4.3).
Let $h$ be a pseudo-Anosov element of $B_{4}$. Then there exists $M_{0}\in\mathbb{N}_{>0}$ such that, for all multiples $M$ of $M_{0}$, we have that $B_{4}/\langle\hskip-1.99168pt\langle h^{M}\rangle\hskip-1.99168pt\rangle$ is a bounded central extension of a HHG, and in particular it is a HHG.
Outline of sections and arguments
We summarise here the techniques that are involved in our arguments. In what follows, we often restrict to central extensions with infinite cyclic kernel for the ease of exposition; in this setting, a quasihomomorphism with image in $\mathbb{Z}$ is simply called a quasimorphism.
In Section 2 we discuss generalities on central extensions and their quotients, boundedness, and quasihomomorphisms. It is worth mentioning Lemma 2.12, which states that, if a quotient extension is bounded, then so is the original extension. Next, we prove the characterisation of bounded quotient extensions, Proposition 2.16. The key observation is that a central extension $1\to\mathbb{Z}\to E\xrightarrow{\pi}G\to 1$ is bounded if and only if it admits a quasimorphism $\phi\colon E\to\mathbb{Z}$ which is the identity on the kernel (see Proposition 2.9). The reader should think of $\phi$ as a “coarse homomorphic” retraction inducing a “coarse” splitting of the extension, as explained below.
Section 3 starts with background material on hierarchically hyperbolic groups (HHG). Roughly, a group $\Gamma$ is a HHG if there exists a collection of hyperbolic spaces $\{\mathcal{C}U\}_{U\in\mathfrak{S}}$, together with coarsely Lipschitz “coordinate projections” $\pi_{U}\colon\Gamma\to\mathcal{C}U$ satisfying various conditions.
We uncover a connection between boundedness and hierarchical hyperbolicity of a central extension. In one direction, Theorem 3.11 shows that a bounded central extension of a HHG is itself a HHG. With an inductive argument, one can reduce to the case where the kernel is $\mathbb{Z}$, which was already settled by [uniform_undistortion, Theorem 5.14]; in turn, the proof of the latter is a refinement of [HRSS_3manifold, Corollary 4.3], which solved the case where the base is hyperbolic. We sketch here the core idea of both arguments, as it enlightens several ideas that appear repeatedly throughout our paper. From a bounded extension one gets an unbounded quasimorphism $\phi\colon E\to\mathbb{Z}$ as above, and the function $E\to\mathbb{Z}\times G$ mapping $e\in E$ to $(\phi(e),\pi(e))$ is a quasi-isometry. In other words, a bounded extension is also quasi-isometrically trivial in the sense of [Gersten_B_is_QIT], though the converse implication is not true in general (see [frigsisto, ascarimilizia] and Remark 3.13). Then the HHG structure for $E$ will have one domain for every domain of $G$, with projection factoring through the quotient map $\pi$, and one quasiline to “detect” the $\mathbb{Z}$-factor. Such quasiline is built out of the quasimorphism $\phi$, using an observation of Abbott, Balasubramanya, and Osin [ABO, Lemma 4.15].
In the opposite direction, Theorem 3.14 proves that, if $E$ is a HHG, then the central extension is bounded, moreover, under the further technical assumption of clean containers, the quotient is a finite-index overgroup of a HHG. To get boundedness we again look for a quasimorphism on $E$ which is the identity on $\mathbb{Z}$. Towards this, one first finds a domain $\mathcal{C}U$ on which the centre acts loxodromically, which must be a quasiline; then the required quasimorphism is the Busemann quasimorphism [Manning_actions_on_hyp, Section 4.1] associated to the action of (a finite-index subgroup of) $E$ on $\mathcal{C}U$, which roughly maps each $e\in E$ to its asymptotic translation length on the quasiline.
We now sketch how to prove hierarchical hyperbolicity of a finite-index subgroup of $G$. Let $\mathcal{C}U_{1},\ldots,\mathcal{C}U_{k}$ be the domains on which $\mathbb{Z}$ acts loxodromically. By taking linear combinations of the associated Busemann quasimorphisms, and then replacing each $\mathcal{C}U_{i}$ with a new quasiline (again using [ABO, Lemma 4.15]), we can modify the HHG structure for $E$ in such a way that $\mathbb{Z}$ acts loxodromically on a single $\mathcal{C}U_{i}$, call it $\mathcal{C}U$, and with uniformly bounded orbits on every other domain. Then the HHG structure of the quotient roughly coincides with what is left of the structure for $E$ after we delete $\mathcal{C}U$.
In the above procedure, one has to restrict to a finite-index subgroup of $E$. This is unavoidable, as in Remark 3.15 we show that, while the $(3,3,3)$ triangle group $G$ is not a HHG by [PS_unbounded], the direct product $\mathbb{Z}\times G$ is. In this sense, the statement of Theorem 1.4 is optimal.
In Section 4 we clarify how the study of hierarchical hyperbolicity of quotients of mapping class groups leads to questions on the boundedness of certain central extensions (Questions 4.1 and 4.2), and discuss possible strategies to tackle them.
Focusing on quotients by the normal closure of a high power of a pseudo-Anosov element, we solve the problem for the 4-strand braid group $B_{4}$ in Theorem 4.3. The key idea is that, if a further quotient extension has hyperbolic base, then it is bounded by [neumannreeves], and hence so is the original quotient extension by Lemma 2.12. Thus the problem becomes to construct a suitable hyperbolic quotient of a five-punctured sphere, which we do using the machinery of rotating families from [dahmani:rotating] and of short HHG from [short_HHG:I].
For surfaces of higher genus, one can maybe adapt the above argument, conditionally to the residual finiteness of certain hyperbolic groups (see Remark 4.5), or try to exploit that the kernel is an increasing union of convex-cocompact subgroups (see Subsection 4.3).
Acknowledgments
We would like to thank Federica Bertolotti, Roberto Frigerio, Shuhei Maruyama, Francesco Milizia, and Davide Spriano for helpful discussions. We are also grateful to Mark Hagen for explaining to us how to cubulate suitable extensions of crystallographic groups. FFF is supported by the Herchel Smith Postdoctoral Fellowship Fund. GM is funded by an EPSRC-DTP PhD studentship.
2. Bounded extensions and quasi(homo)morphisms
In this section we discuss generalities on central extensions and bounded cohomology. In particular, we state a characterisation of boundedness in terms of quasimorphisms (Proposition 2.9). The main result in this section is Proposition 2.16, which gives equivalent characterisations for a quotient central extension to be bounded.
Notation 2.1.
Throughout this section, we use $\sim$ to denote equality up to a uniformly bounded error.
Definition 2.2.
In this paper, cocycle will always refer to inhomogeneous $2$-cocycle, namely a map $\omega\colon G^{2}\to K$ such that
$\delta\omega(g_{1},g_{2},g_{3})=\omega(g_{2},g_{3})-\omega(g_{1}g_{2},g_{3})+\omega(g_{1},g_{2}g_{3})-\omega(g_{1},g_{2})=0$
for all $g_{1},g_{2},g_{3}\in G$.
Let us briefly recall the dictionary between second cohomology and central extensions, referring the reader to [brown, Chapter IV.3] for details. Given a central extension $1\to K\to E\to G\to 1$, we pick a section $\sigma\colon G\to E$ that is normalised, meaning that $\sigma(1)=1$. We identify $\omega(g,h)=\sigma(g)\sigma(h)\sigma(gh)^{-1}$ with an element of $K$, and then $\omega\colon G^{2}\to K$ is a cocycle that is normalised, meaning that $\omega(1,g)=\omega(g,1)=0$ for all $g\in G$. This defines a class $\omega]\in H^{2}(G;K)$, independent of the choice of $\sigma$, which is called the Euler class of the extension and we also denote by $[E]$. Conversely, given a normalised cocycle $\omega\colon G^{2}\to K$ we define a group $E$ with underlying set $K\times G$ and product
$(k_{1},g_{1})\cdot(k_{2},g_{2})=(k_{1}+k_{2}+\omega(g_{1},g_{2}),g_{1}g_{2}).$
Then $E$ is a central extension as above, and the map $g\to(0,g)$ is a normalised section. Equality in cohomology corresponds to equivalence of central extensions.
Definition 2.3.
Let $(K,\|\cdot\|)$ be an Abelian group endowed with a norm. For another group $G$, a class $\alpha\in H^{2}(G;K)$ is bounded if it belongs to the image of the comparison map $H^{2}_{b}(G;K)\to H^{2}(G;K)$. More explicitly, $\alpha$ is bounded if there exists a cocycle $\omega:G^{2}\to K$ such that $\alpha=[\omega]$, and such that $\|\omega(\cdot,\cdot)\|\colon G^{2}\to\mathbb{Z}$ is uniformly bounded. A central extension $1\to K\to E\to G\to 1$ is bounded if the corresponding Euler class $E]\in H^{2}(G;K)$ is bounded.
Definition 2.4.
Let $(K,\|\cdot\|)$ be as above, and let $E$ be another group. A map $\chi\colon E\to K$ is a quasihomomorphism if there exists $D(\chi)\geq 0$, called the defect of $\chi$, such that, for every $e_{1},e_{2}^{\prime}\in E$,
$\|\chi(e_{1})+\chi(e_{2})-\chi(e_{1}e_{2})\|\leq D(\chi).$
A quasihomomorphism is homogeneous if it restricts to a homomorphism on every cyclic subgroup. When $K=\mathbb{Z}$ or $\mathbb{R}$ with the Euclidean norm, we say quasimorphism instead of quasihomomorphism.
Remark 2.5 (Homogeneisation).
Given a quasimorphism $\phi\colon E\to\mathbb{R}$, for every $g\in E$ the limit $\phi_{h}(g)=\lim_{n\to\infty}\frac{\phi(g^{n})}{n}$ always exists. The map $\phi_{h}\colon E\to\mathbb{R}$ is a homogeneous quasimorphism; moreover, both $\|\phi-\phi_{h}\|$ and the defect of $\phi_{h}$ are bounded in terms of the defect of $\phi$ [calegari, Lemma 2.21].
Real-valued quasimorphisms will only appear in the course of proofs about quasimorphisms with values in $\mathbb{Z}$. All of our statements will only involve the case in which $K$ is a finitely generated Abelian group endowed with a word norm. In this case, a subset of $K$ is bounded if and only if it is finite, so the notions above are independent of the choice of a norm. This allows to generalise the notion of quasihomomorphism to maps taking values in any discrete group, following Fujiwara–Kapovich [fujikapo].
Definition 2.6.
A map $\chi\colon G\to H$ between groups is a quasihomomorphism if there exist a finite set $F\subset H$ such that, for every $g_{1},g_{2}\in G$, the difference between $\chi(g_{1}g_{2})$ and $\chi(g_{1})\chi(g_{2})$ lies in $F$ (the way in which one takes the difference does not matter by [heuer, Proposition 2.3]).
We record here an example of a quasimorphism that we shall use repeatedly throughout the paper.
Example 2.7 (Busemann quasimorphism, see e.g. [Manning_actions_on_hyp, Section 4.1]).
Let $E$ be a group acting on a $\delta$-hyperbolic metric space $X$, with Gromov boundary $\partial X$, and suppose the action fixes an ideal point $p\in\partial X$. Given a sequence $\{x_{n}\}\in X$ converging to $p$, the map $\phi_{\{x\}}\colon E\to\mathbb{R}$ defined by
$\phi_{\{x\}}(g)=\limsup_{n\to\infty}\mathrm{d}(gx_{0},x_{n})-\mathrm{d}(x_{0},x_{n}),$
is a quasimorphism of defect at most $16\delta$. Then the Busemann quasimorphism $\phi$ is the homogeneisation of $\phi_{\{x\}}$, as in Remark 2.5. One can check that $\phi$ does not depend on the choice of the sequence; moreover, an element $g\in E$ acts loxodromically on $X$ if and only if $\phi(g)\neq 0$. Note that, by construction, for every $g\in E$ one has that $|\phi(g)|\leq\mathrm{d}(x_{0},gx_{0})$; conversely, if $X$ is a quasiline, then there exists a constant $L$ such that $\mathrm{d}(x_{0},gx_{0})\leq|\mathrm{d}(gx_{0},x_{n})-\mathrm{d}(x_{0},x_{n})|+L$, for all sufficiently large $n$ (if $\{x_{0},gx_{0},x_{n}\}=\{a,b,c\}$ and $b$ is “between” $a$ and $c$, one can see this by considering the distance between $b$ and a geodesic $[a,c]$). Taking the limsup, one gets that $\mathrm{d}(x_{0},gx_{0})$ and $|\phi(g)|$ are within uniform distance.
We now turn to the characterisation of when a central extension is bounded, in terms of the existence of certain quasihomomorphisms.
Lemma 2.8.
Let $1\to K\to E\to G\to 1$ be a central extension.
  • Every bounded cocycle $\omega\colon G^{2}\to K$ is cohomologous to a normalised bounded cocycle.
  • If there exists a quasihomomorphic section $\sigma\colon G\to E$, then there exists a normalised quasihomomorphic section.
Proof.
For the first bullet, we have
$0=\delta\omega(g^{-1},g,1)=\omega(g,1)-\omega(1,1)+\omega(g^{-1},g)-\omega(g^{-1},g).$
Therefore $\omega(g,1)=\omega(1,1)$, and similarly $\omega(1,g)=\omega(1,1)$ for all $g\in G$. Let $b\colon G\to K$ be the constant function at $\omega(1,1)$. Then $\delta b(g,h)=\omega(1,1)$ for all $g,h\in G$. Setting $\omega^{\prime}=\omega-\delta b$, we see that $\omega^{\prime}$ is a normalised bounded cocycle cohomologous to $\omega$.
For the second bullet, let $\sigma^{\prime}(g)=\sigma(g)$ for all $g\neq 1$, and $\sigma^{\prime}(1)=1$. We need to check that $\sigma^{\prime}(g_{1})\sigma^{\prime}(g_{2})\sigma(g_{1}g_{2})^{-1}$ takes finitely many values over $g_{1},g_{2}\in G$. If $g_{1},g_{2},g_{1}g_{2}\neq 1$, then this follows from $\sigma$ being a quasihomomorphism. If one of $g_{1},g_{2}$ is equal to $1$, then the equation above is just equal to $1$. Finally, if $g_{1}=g\neq 1$ and $g_{2}=g^{-1}$, then the equation above is equal to
$\sigma(g)\sigma(g^{-1})=\big(\sigma(g)\sigma(g^{-1})\sigma(1)^{-1}\big)\sigma(1),$
which again takes finitely many values because $\sigma$ is a quasihomomorphism. ∎
Proposition 2.9.
Let $1\to K\to E\xrightarrow{\pi}G\to 1$ be a central extension with finitely generated kernel. The following are equivalent:
  1. (1)
    $[E]$ is bounded, i.e. $[E]$ is represented by a (normalised) bounded cocycle.
  2. (2)
    There exists a quasihomomorphism $\chi\colon E\to K$ such that $\chi|_{K}$ is the identity.
  3. (3)
    There exists a (normalised) section $s\colon G\to E$ which is a quasihomomorphism.
Shortly after this paper was uploaded, Tao and Wan independently proved Proposition 2.9 with roughly the same methods (see [tao2025properactionsfiniteproducts, Lemma 5.6]).
Proof.
By Lemma 2.8, we may assume that all cocycles and sections involved are normalised. Up to isomorphism of central extensions, $E$ has underlying set $K\times G$, and product
$(k_{1},g_{1})\cdot(k_{2},g_{2})=(k_{1}+k_{2}+\omega(g_{1},g_{2}),g_{1}g_{2});$
where $\omega\colon G^{2}\to K$ is a normalised cocycle representing the central extension.
[…]
$\chi((k_{1},g_{1})\cdot(k_{2},g_{2}))=k_{1}+k_{2}+\omega(g_{1},g_{2})=\chi(k_{1},g_{1})+\chi(k_{2},g_{2})+\omega(g_{1},g_{2}).$
Thus the defect of $\chi$ is bounded by the norm of $\omega$, and so it is bounded, i.e. $\chi$ is a quasihomomorphism.
[…]
$(\omega-\delta b)(g_{1},g_{2})$ $=\omega(g_{1},g_{2})+b(g_{1}g_{2})-b(g_{1})-b(g_{2})$
$=\chi(\omega(g_{1},g_{2}),1_{G})+\chi(0,g_{1}g_{2})-\chi(0,g_{1})-\chi(0,g_{2})$
$\sim\chi(\omega(g_{1},g_{2}),g_{1}g_{2})-\chi(0,g_{1})-\chi(0,g_{2})$
$=\chi((0,g_{1})\cdot(0,g_{2}))-\chi(0,g_{1})-\chi(0,g_{2})\sim 0.$
Therefore $\omega-\delta b$ is bounded, and it is a cocycle cohomologous to $\omega$. This shows that the extension is bounded.
[…]
Corollary 2.10.
Let $1\to K\to E\xrightarrow{\pi}G\to 1$ be a bounded central extension with finitely generated kernel. Let $L<K$ be a subgroup. Then the central extensions
$1\to L\to E\to E/L\to 1$
and
$1\to K/L\to E/L\to G\to 1$
are also bounded.
Proof.
We use the characterisation in the second item of Proposition 2.9. Let $\chi\colon E\to K$ be a quasihomomorphism such that $\chi|_{K}$ is the identity. Then $\chi|_{L}$ is the identity, which proves that the $L$-central extension is bounded. Composing $\chi$ with the projection $K\to K/L$, we get a quasihomomorphism $E/L\to K/L$ that is the identity on $K/L$, whcih proves that the $K/L$-central extension is bounded. ∎
Given a central extension $1\to K\to E\xrightarrow{\pi}G\to 1$, let $N\leq E$ be a normal subgroup that intersects $K$ trivially, so that $\pi\colon N\to\pi(N)$ is an isomorphism. Then there is a diagram
[…]
where $\overline{E}=E/N$ and $\overline{G}=G/\pi(N)$, and both rows are central extension. We use the letter $p$ to denote both quotients $E\to\overline{E}$ and $G\to\overline{G}$ by an abuse of notation. We say that the bottom row is a quotient central extension of the top row. The main focus of the paper is the following question (see Problem 1.1):
Question 2.11.
Under which conditions is a quotient central extension bounded?
A first easy necessary condition is given by the following observation:
Lemma 2.12.
With the notation of Diagram (1), the pullback $p^{*}\colon H^{2}(\overline{G};K)\to H^{2}(G;K)$ sends $\overline{E}]$ to $[E]$. In particular, if $\overline{E}]$ is bounded, then $[E]$ is bounded.
Proof.
We choose a normalised section $\overline{\sigma}\colon\overline{G}\to\overline{E}$. We also choose an injective map $\tau\colon\overline{G}\to E$ such that $p\tau=\overline{\sigma}$ and $\tau(1_{\overline{G}})=1_{E}$. This way Diagram (1) is enriched as follows:
[…]
Now $\pi\tau$ is a section for $p\colon G\to\overline{G}$, so every element of $G$ can be written uniquely as $g=\pi(n)\cdot\pi\tau(p(g))$ for some $n\in N$. We define $\sigma(g)=n\cdot\tau(p(g))$. First, note that $\sigma$ is indeed a normalised section: $\sigma(1_{G})=1_{E}$, and if $g=\pi(n)\pi\tau(p(g))$, then
$\pi\sigma(g)=\pi(n\cdot\tau(p(g)))=\pi(n)\cdot\pi\tau(p(g))=g.$
Secondly, we claim that $p\sigma=\overline{\sigma}p$, indeed
$p\sigma(g)=p(n\cdot\tau(p(g)))=p\tau(p(g))=\overline{\sigma}(p(g)),$
which implies that the projection $p\colon E\to\overline{E}$ sends
$\sigma(g)\sigma(h)\sigma(gh)^{-1}\mapsto\overline{\sigma}(p(g))\overline{\sigma}(p(h))\overline{\sigma}(p(gh))^{-1}.$
This shows that the cocycle defined by $\sigma$ is indeed the pullback of the cocycle defined by $\overline{\sigma}$, and concludes the proof. ∎
This condition is however not sufficient, as the following examples show:
Example 2.13.
Consider the group $Z\times F_{2}$, where $Z=\langle z\rangle$ is infinite cyclic, and $F_{2}=\langle x,y\rangle$ is free of rank $2$. Let $N$ be the normal closure of $z^{-1}[x,y]$, which intersects $Z$ trivially. The quotient $(Z\times F_{2})/N$ is the Heisenberg group $H_{3}$, and the quotient $F_{2}/\pi(N)$ is the free Abelian group $\mathbb{Z}^{2}$. So we have a map of central extensions:
[…]
The top one splits, so it is trivial and in particularly bounded. The bottom one is not trivial, and in fact it maps to a generator of $H^{2}(\mathbb{Z}^{2};\mathbb{R})\cong\mathbb{R}$ under the change of coefficients map. Because $\mathbb{Z}^{2}$ is amenable, $H^{2}_{b}(\mathbb{Z}^{2};\mathbb{R})$ vanishes [frigerio, Chapter 3], and so the bottom central extension cannot be bounded.
Example 2.14.
More generally, we claim that every central extension can be expressed as a quotient of a trivial (thus bounded) central extension of a free group. Let $1\to K\to E\to G\to 1$ be a central extension. Let $F$ be a free group and $p\colon F\to G$ a presentation, which we lift to a homomorphism $\tilde{p}\colon F\to E$. Define $P\colon K\times F\to E$ to be the product of $\tilde{p}$ and the inclusion $K\to E$. An element $(k,w)\in K\times F$ belongs to the kernel of $P$ if and only if $\tilde{p}(w)=k$; in particular $K$ intersects this kernel trivially. So we indeed have a quotient central extension:
[…]
We can restate Question 2.11 in terms of extendability of a certain quasihomomorphism. We start with an easy lemma, which is a variation of a well-known fact regarding real-valued homogeneous quasimorphisms.
Lemma 2.15.
Let $N\leq G$ be a normal subgroup, and denote by $p\colon G\to G/N$ the quotient. Let $\chi\colon G\to K$ be a quasihomomorphism. If $\chi|_{N}\equiv 0$, then there exists a quasihomomorphism $\phi\colon G/N\to K$ such that $\chi$ is at a bounded distance from $\phi p$. Moreover, if $H\leq G$ is a subgroup such that $\chi|_{H}$ is a homomorphism, then $\phi$ can be chosen so that $\phi p|_{H}=\chi|_{H}$.
Proof.
Let $K=\mathbb{Z}^{k}\times T$, where $k\geq 0$ and $T$ is a finite abelian group. Accordingly, write $\chi=(\chi^{1},\ldots,\chi^{k},\chi^{T})$, and note that $\chi^{1},\ldots,\chi^{k}$ are quasimorphisms (while $\chi^{T}$ can be any function $G\to T$). Recall that every quasimorphism $\chi^{i}\colon G\to\mathbb{R}$ is at a bounded distance from a unique homogeneous quasimorphism $\chi^{i}_{h}$ (Remark 2.5). The hypothesis, and uniqueness of homogeneous representatives, implies that $\chi^{i}_{h}|_{N}\equiv 0$ for all $1\leq i\leq k$. Therefore, by [calegari, Remark 2.90] there exist homogeneous quasimorphisms $\phi^{i}_{h}\colon G/N\to\mathbb{R}$ such that $\chi^{i}_{h}=\phi^{i}_{h}p$ for all $1\leq i\leq k$. Setting $\phi^{i}$ to be the integral part of $\phi^{i}_{h}$, we obtain a quasimorphism $\phi^{i}\colon G/N\to\mathbb{Z}$ such that $\phi^{i}p\sim\phi^{i}_{h}p=\chi^{i}_{h}\sim\chi^{i}$. Setting $\phi=(\phi^{1},\ldots,\phi^{k},\phi^{T})$, where $\phi^{T}\colon G/N\to T$ is any function, we obtain that $\chi$ is at a bounded distance from $\phi p$. Notice that $\phi$ is indeed a quasihomomorphism, as its entries are quasimorphisms.
Suppose that $\chi|_{H}$ is a homomorphism. Then by uniqueness of homogeneous representatives, $\chi^{i}|_{H}=\chi^{i}_{h}|_{H}=\phi^{i}_{h}p|_{H}$ for all $i$. In particular, $\phi^{i}_{h}$ takes integer values on $p(H)$, and so $\phi^{i}|_{p(H)}=\phi^{i}_{h}|_{p(H)}$. Moreover, because $\chi|_{N}\equiv 0$, there exists a map $\phi^{T}\colon G/N\to T$ such that $\phi^{T}p|_{H}=\chi^{T}|_{H}$. With this choice of $\phi^{T}$, we have $\phi p|_{H}=\chi|_{H}$, as promised. ∎
Back to the setting of Diagram (1), we now provide the characterisation of boundedness of central extensions in terms of extendable quasihomomorphisms. Recall from Lemma 2.12 that, in order for the quotient central extension to be bounded, the original central extension needs to be bounded, so we assume this throughout.
Proposition 2.16.
With the notation of Diagram (1), suppose that $[E]$ is bounded, and let $\chi\colon E\to K$ be a quasihomomorphism such that $\chi|_{K}$ is the identity, provided by Proposition 2.9. Then:
  • The map $\chi\pi^{-1}\colon\pi(N)\to K$ is a well-defined quasihomomorphism, and it is almost invariant under conjugacy by $G$.
  • If $\omega$ is a normalised bounded cocycle representing $[E]$, then $\chi$ can be chosen so that $\delta(\chi\pi^{-1})=\omega|_{\pi(N)}$.
  • $\overline{E}]$ is bounded if and only if $\chi\pi^{-1}$ can be extended to a quasihomomorphism $G\to K$.
Given a normal subgroup $\Lambda\leq\Gamma$, we say that a quasihomomorphism $\phi\colon\Lambda\to K$ is almost invariant under conjugacy by $\Gamma$ if there exists a constant $C>0$ such that
$\|\phi(\gamma\lambda\gamma^{-1})-\phi(\lambda)\|<C$
for all $\gamma\in\Gamma,\lambda\in\Lambda$. Note that quasihomomorphisms on $\Gamma$ are almost invariant under conjugacy by $\Gamma$, so this is a necessary condition for a quasihomomorphism on $\Lambda$ to be extendable, which will come up in the next result.
Proof of Proposition 2.16.
$\chi\pi^{-1}$ is well-defined because $\pi|_{N}$ is an isomorphism. Because $\chi$ is defined on $E$, it is almost invariant under conjugacy by $E$, and then it follows that $\chi\pi^{-1}$ is almost invariant under conjugacy by $G$.
Moving to the second bullet, from the proof of Proposition 2.9 we see that if $\omega$ is a normalised bounded cocycle representing $E$, seen as the Cartesian product $K\times G$ with the product twisted by $\omega$, then $\chi$ can be chosen to be the projection onto the first factor, while $\pi$ is the projection onto the second factor. Now every element $g\in N$ can be written as $(b(g),\pi(g))$, and it follows that $\chi\pi^{-1}\colon\pi(N)\to K$ is just the map $\pi(g)\mapsto b(g)$. Because $N$ is a group and $\pi$ is injective, for all $g_{1},g_{2}\in N$ it holds
$(b(g_{1}g_{2}),\pi(g_{1}g_{2}))$ $=(b(g_{1}),\pi(g_{1}))\cdot(b(g_{2}),\pi(g_{2}))$
$=(b(g_{1})+b(g_{2})+\omega(\pi(g_{1}),\pi(g_{2})),\pi(g_{1}g_{2}));$
which shows that $\delta(\chi\pi^{-1})=\omega$ as cocycles $\pi(N)^{2}\to K$.
We finally prove the equivalence in the third bullet. Suppose that $\chi\pi^{-1}$ extends to a quasihomomorphism $\phi\colon G\to K$. Consider the quasihomomorphism $\chi-\phi\pi\colon E\to K$. Notice that $\chi-\phi\pi$ restricts to the zero map on $N$ and to the identity on $K$. Therefore, Lemma 2.15 yields a quasihomomorphism $\psi\colon\overline{E}\to K$ such that $\psi p$ is at a bounded distance from $\chi-\phi\pi$, and restricts to the identity on $K$. Thus $\overline{E}]$ is bounded, by the second item of Proposition 2.9.
Conversely, suppose that $\psi\colon\overline{E}\to K$ is a quasihomomorphism that restricts to the identity on $K$. Then $\psi p\colon E\to K$ vanishes on $N$ and is still the identity on $K$. Therefore $\chi-\psi p\colon E\to K$ vanishes on $K$, and coincides with $\chi$ on $N$. By Lemma 2.15, there exists a quasihomomorphism $\phi\colon G\to K$ such that $\phi\pi\colon E\to K$ is at a bounded distance from $\chi-\psi p$. So, up to changing $\phi$ by a bounded amount on $\pi(N)$, it is an extension of $\chi\pi^{-1}$. ∎
Remark 2.17.
Shuhei Maruyama pointed out to us that both Propositions 2.9 and 2.16 can also be proved by diagram chasing, using the exact sequences from [extendable:main].
In view of Proposition 2.16, the problem of whether the quotient central extension $\overline{E}]$ is bounded is equivalent to the problem of whether a specific quasihomomorphism $\pi(N)\to K$ extends to $G$. This problem is non-trivial. On the one hand, the easy obstruction for extendability, that is almost invariance under conjugacy by $G$, is excluded by the the first bullet of Proposition 2.16. On the other hand, the almost invariance under conjugacy is not sufficient for a quasihomomorphism to be extendable in general, as we have seen in Example 2.13. The problem of extendability has been studied at length in recent years, and produced various examples of non-extendable quasimorphisms, and criteria for extendability: see [extendable:survey] for a survey. In particular, we recall the following criteria:
Proposition 2.18 ([extendable:bavard, Proposition 1.6] and [extendable:main, Theorem 1.9]).
Let $1\to N\to G\to G/N\to 1$ be a short exact sequence of groups. Then any of the following conditions guarantees the extendability of every quasihomomorphism $N\to K$ that is almost invariant under conjugacy by $G$.
  1. (1)
    The extension virtually splits.
  2. (2)
    $H^{3}_{b}(G/N;\mathbb{R})=H^{2}_{b}(G/N;\mathbb{R})=H^{2}(G/N;\mathbb{R})=0$.
In another direction, very far from the case of a normal subgroup, we have:
Proposition 2.19 ([Hull_Osin], see also [FPS]).
Let $H\leq G$ be a hyperbolically embedded subgroup. Then every quasihomomorphism $H\to K$ extends to a quasihomomorphism $G\to K$.
For both propositions, the original statement is about real-valued quasimorphisms, however these statements follow by splitting $K$ as $\mathbb{Z}^{k}\times T$, choosing arbitrary extensions for the $T$ part, and dealing with the $\mathbb{Z}$ factors by taking the integral part and modifying by a bounded amount.
Remark 2.20.
By Proposition 2.16, the quasihomomorphism one has to extend is always such that its boundary is extendable. This shows a stark difference from the problem of extendability of bounded cohomology classes and extendability of quasimorphisms [extendable:survey, End of Section 1.3].
3. Hierarchical hyperbolicity of central extensions
In this Section we explore the connections between boundedness and hierarchical hyperbolicity for central extensions. We first recall the notion of hierarchically hyperbolic spaces and groups.
Definition 3.1 (Hierarchically hyperbolic space).
Let $\delta>0$ and $\mathcal{X}$ be a $(\delta,\delta)$–quasigeodesic space. A hierarchically hyperbolic space (HHS) structure with constant $\delta$ for $\mathcal{X}$ is the data of an index set $\mathfrak{S}$ and a set $\{\mathcal{C}W\colon W\in\mathfrak{S}\}$ of $\delta$-hyperbolic spaces $(\mathcal{C}W,\mathrm{d}_{W})$ such that the following axioms are satisfied.
  1. (1)
    (Projections.) For each $W\in\mathfrak{S}$, there exists a projection $\pi_{W}\colon\mathcal{X}\rightarrow 2^{\mathcal{C}W}$ that is a $(\delta,\delta)$–coarsely Lipschitz, $\delta$–coarsely onto, $\delta$–coarse map.
  2. (2)
[…]
  3. (3)
[…]
  4. (4)
[…]
  5. (5)
    (Containers.) For each $W\in\mathfrak{S}$ and $U\in\mathfrak{S}_{W}$ with $\mathfrak{S}_{W}\cap\mathfrak{S}_{U}^{\perp}\neq\emptyset$, there exists $Q\in\mathfrak{S}_{W}-\{W\}$ such that $V\sqsubseteq Q$ whenever $V\in\mathfrak{S}_{W}\cap\mathfrak{S}_{U}^{\perp}$. We call $Q$ the container of $U$ in $W$.
  6. (6)
    (Transversality.) If $V,W\in\mathfrak{S}$ are not orthogonal and neither is nested in the other, then we say $V$ and $W$ are transverse, denoted $V\pitchfork W$. Moreover, for all $V,W\in\mathfrak{S}$ with $V\pitchfork W$, there are non-empty sets $\rho^{V}_{W}\subseteq\mathcal{C}W$ and $\rho^{W}_{V}\subseteq\mathcal{C}V$, each of diameter at most $\delta$.
  7. (7)
    (Consistency.) For all $x\in\mathcal{X}$ and $U,V,W\in\mathfrak{S}$:
    • if $V\pitchfork W$, then $\min\left\{\mathrm{d}_{W}(\pi_{W}(x),\rho^{V}_{W}),\mathrm{d}_{V}(\pi_{V}(x),\rho^{W}_{V})\right\}\leq\delta$,
    • if $U\sqsubseteq V$ and either $V\sqsubsetneq W$, or $V\pitchfork W$ and $W\not\perp U$, then $\mathrm{d}_{W}(\rho^{U}_{W},\rho^{V}_{W})\leq\delta$.
  8. (8)
    (Bounded geodesic image (BGI).) For all $V,W\in\mathfrak{S}$ and for all $x,y\in\mathcal{X}$, if $V\sqsubsetneq W$ and $\mathrm{d}_{V}(\pi_{V}(x),\pi_{V}(y))\geq\delta$, then every $\mathcal{C}W$–geodesic from $\pi_{W}(x)$ to $\pi_{W}(y)$ must intersect $\mathcal{N}_{\delta}(\rho_{W}^{V})$.
  9. (9)
    (Large links.) For all $W\in\mathfrak{S}$ and $x,y\in\mathcal{X}$, there exists a collection $\{V_{1},\dots,V_{m}\}\subseteq\mathfrak{S}_{W}-\{W\}$ such that $m\leq\delta\mathrm{d}_{W}(\pi_{W}(x),\pi_{W}(y))+\delta$, and for all $U\in\mathfrak{S}_{W}-\{W\}$, either $U\sqsubseteq V_{i}$ for some $i$, or $\mathrm{d}_{U}(\pi_{U}(x),\pi_{U}(y))\leq\delta$.
  10. (10)
    (Partial realization.) If $\{V_{i}\}$ is a finite collection of pairwise orthogonal elements of $\mathfrak{S}$ and $p_{i}\in\mathcal{C}V_{i}$ for each $i$, then there exists $x\in\mathcal{X}$ realising the tuple $(p_{i})$, meaning that, for every $i$ and every $W\in\mathfrak{S}$:
    • $\mathrm{d}_{V_{i}}(\pi_{V_{i}}(x),p_{i})\leq\delta$;
    • if $V_{i}\sqsubsetneq W$ or $W\pitchfork V_{i}$ then $\mathrm{d}_{W}(\pi_{W}(x),\rho^{V_{i}}_{W})\leq\delta$.
  11. (11)
    (Uniqueness.) There exists a function $\theta\colon[0,\infty)\to[0,\infty)$ so that for all $r\geq 0$, if $x,y\in\mathcal{X}$ and $\mathrm{d}_{\mathcal{X}}(x,y)\geq\theta(r)$, then there exists $W\in\mathfrak{S}$ such that $\mathrm{d}_{W}(\pi_{W}(x),\pi_{W}(y))\geq r$.
We use $\mathfrak{S}$ to denote the HHS structure. We call an element $U\in\mathfrak{S}$ a domain, the associated space $\mathcal{C}U$ its coordinate space, and call the maps $\rho_{W}^{V}$ the relative projections from $V$ to $W$. The quantity $\delta$ is called a hierarchy constant for $\mathfrak{S}$. We often suppress reference to the projection maps, so for every $x,y\in\mathcal{X}$ and $U\in\mathfrak{S}$ we write $\mathrm{d}_{U}(x,y)$ to mean $\mathrm{d}_{U}(\pi_{U}(x),\pi_{U}(y))$.
Definition 3.2 (Hierarchically hyperbolic group).
A finitely generated group $G$ is a hierarchically hyperbolic group (HHG) if the following hold.
  1. (i)
    $G$ acts metrically properly and coboundedly on a space $\mathcal{X}$ admitting a HHS structure $\mathfrak{S}$.
  2. (ii)
[…]
  3. (iii)
    For each $W\in\mathfrak{S}$ and $g\in G$, there exists an isometry $g_{W}\colon\mathcal{C}W\to\mathcal{C}(gW)$ satisfying the following for all $V,W\in\mathfrak{S}$ and $g,h\in G$.
    • The maps $(gh)_{W}\colon\mathcal{C}W\to\mathcal{C}(ghW)$ and $g_{hW}\circ h_{W}\colon\mathcal{C}W\to\mathcal{C}(hW)$ coincide.
    • For each $x\in X$, $g_{W}(\pi_{W}(x))=\pi_{gW}(g\cdot x)$ in $\mathcal{C}(gW)$.
    • If $V\pitchfork W$ or $V\sqsubsetneq W$, then $g_{W}(\rho^{V}_{W})=\rho^{gV}_{gW}$ in $\mathcal{C}(gW)$.
We often drop the indices and denote each $g_{W}$ simply by $g$. When the underlying HHS is not relevant, we denote a HHG by $(G,\mathfrak{S})$.
Remark 3.3 (Moral compass).
When reading the definitions above, the uninitiated reader should keep in mind the motivating example of a HHG, which is the mapping class group of an orientable, finite-type surface. In this context, $\mathcal{X}$ is the marking complex from [masurminsky2, Section 2.5]; the elements of $\mathfrak{S}$ are isotopy classes of subsurfaces, with nesting given by inclusion and orthogonality corresponding to disjointness (both up to isotopy); finally, the coordinate space associated to a subsurface is the corresponding curve graph, onto which $\mathcal{X}$ maps via the subsurface projection. The various axioms from Definition 3.1 are abstractions of the properties of curve graphs and clean markings from [masurminsky1, masurminsky2].
3.1. Variations on the axioms
Remark 3.4 (Normalisation).
For experts, Definition 3.1 is that of a normalised HHS, since the original definition [HHS_II] only required the coordinate projections to have uniformly quasiconvex images. However it is always possible to normalise a HHS structure, by restricting each coordinate space $\mathcal{C}U$ to a neighbourhood of $\pi_{U}(\mathcal{X})$ (see e.g. [HHS_II, Remark 1.3]).
The following is a strengthening of Axiom (5):
Definition 3.5.
A hierarchically hyperbolic space $(\mathcal{X},\mathfrak{S})$ has clean containers if for each $W\in\mathfrak{S}$ and $U\in\mathfrak{S}_{W}$ with $\mathfrak{S}_{W}\cap\mathfrak{S}_{U}^{\perp}\neq\emptyset$, there exists $Q\in\mathfrak{S}_{W}-\{W\}$ such that $V\sqsubseteq Q$ whenever $V\in\mathfrak{S}_{W}\cap\mathfrak{S}_{U}^{\perp}$, and $Q\bot U$.
Remark 3.6 (Bounded domain dichotomy).
A HHG acts cofinitely on the set of domains, and two domains in the same orbit have isometric coordinate spaces. Therefore, up to enlarging $\delta$, we can and will assume that every $U\in\mathfrak{S}$ is either unbounded (meaning that $\mathcal{C}U$ is) or $\mathrm{diam}\mathcal{C}U\leq\delta$.
Remark 3.7 (Passing up).
The large links Axiom (9) can be replaced by the following weaker requirement:
  1. (9’)
    (Passing up.) For every $t>0$, there exists an integer $P=P(t)>0$ such that if $V\in\mathfrak{S}$ and $x,y\in\mathcal{X}$ satisfy $\mathrm{d}_{U_{i}}(x,y)>\delta$ for a collection of domains $\{U_{i}\}_{i=1}^{P}$ with $U_{i}\in\mathfrak{S}_{V}$, then there exists $W\in\mathfrak{S}_{V}$ containing some $U_{i}$, and such that $\mathrm{d}_{W}(x,y)>t$. We call $P\colon(0,\infty)\to(0,\infty)$ the passing up function.
It was shown in [HHS_II, Lemma 2.5] that every HHS satisfies the Passing up axiom; conversely, in [Durham_cubinf, Section 4.8] it is argued that the passing up axiom, together with BGI and normalisation, imply the large links axiom.
Remark 3.8.
Notice that, when one has to verify the Passing up axiom, it is enough to define $P(t)$ only for $t\geq\delta$, because then one can set
[…]
In particular, if the candidate collection of domains has the bounded domain dichotomy, one can further require the domains $\{U_{i}\}_{i=1}^{P}$ and $W$ to be unbounded.
3.2. Action of the centre
Here we prove several results about the action of the centre of a HHG on the structure. Firstly, we recall a definition:
Definition 3.9 (Big domains [DHS]).
Let $(\mathcal{X},\mathfrak{S})$ be a HHG structure for $G$, fix a basepoint $x\in\mathcal{X}$, and let $z\in G$. The bigset of $z$ is
$\operatorname{Big}\left(z\right)=\{U\in\mathfrak{S}\,|\,\mathrm{diam}\pi_{U}(\langle z\rangle\cdot x)=\infty\},$
that is, the set of coordinate spaces “witnessing” the infinite order of $z$.
It is clear that $\operatorname{Big}\left(z\right)$ does not depend on the choice of the basepoint. We have that $z$ has infinite order if and only if $\operatorname{Big}\left(z\right)\neq\emptyset$ [DHS, Proposition 6.4], and that $\operatorname{Big}\left(z\right)$ is a finite collection of pairwise orthogonal domains (this follows from combining [DHS, Lemma 6.7] with [HHS_II, Lemma 2.1]).
Lemma 3.10.
Let $(\mathcal{X},\mathfrak{S})$ be a HHG structure for a group $E$, and let $K$ be finitely generated and contained in the centre of $E$. Then:
  1. (1)
    $K$ fixes every unbounded domain.
  2. (2)
    Let $\{z_{1},\ldots,z_{n}\}$ be a basis for $K$. Then there exists $C>0$ such that, for every $x\in\mathcal{X}$ and every $W\in\mathfrak{S}-\bigcup_{i=1}^{n}\operatorname{Big}\left(z_{i}\right)$, $\mathrm{diam}\pi_{W}(K\cdot x)<C$.
Proof.
(1) Let $\delta$ be a HHG constant for the structure. We first argue that the $K$-action on $\mathfrak{S}$ fixes every unbounded domain. Fix $x_{0}\in\mathcal{X}$ and $k\in K-\{1\}$, and let $W$ be any unbounded domain. For every $g,h\in E$,
$\mathrm{d}_{kW}(hx_{0},ghx_{0})=\mathrm{d}_{W}(k^{-1}hx_{0},k^{-1}ghx_{0});$
then the triangle inequality, plus the fact that $k$ is central in $E$, yield that
$\left|\mathrm{d}_{W}(k^{-1}hx_{0},k^{-1}ghx_{0})-\mathrm{d}_{W}(hx_{0},ghx_{0})\right|$ $\leq\mathrm{d}_{W}(hx_{0},k^{-1}hx_{0})+\mathrm{d}_{W}(ghx_{0},k^{-1}ghx_{0})$
$=\mathrm{d}_{h^{-1}W}(x_{0},k^{-1}x_{0})+\mathrm{d}_{(gh)^{-1}W}(x_{0},k^{-1}x_{0})$
$\leq 2\sup_{V\in\mathfrak{S}}\mathrm{d}_{V}(x_{0},k^{-1}x_{0}).$
The quantity $M\coloneq 2\sup_{V\in\mathfrak{S}}\mathrm{d}_{V}(x_{0},k^{-1}x_{0})$ is bounded in terms of $\mathrm{d}_{\mathcal{X}}(x_{0},k^{-1}x_{0})$, since coordinate projections are uniformly coarsely Lipschitz. We just proved that projection distances in $W$ and in $kW$ must agree up to an error of $M$.
Now we show that $kW=W$, by excluding every other possibility.
  • If, say, $W\sqsubsetneq kW$ and $k$ has infinite order, then we would get an infinite chain of properly nested domains $W\sqsubsetneq kW\sqsubsetneq k^{2}W\ldots$, contradicting finite complexity; if instead $k$ has order $n<\infty$, then the above chain would eventually yield that $W\sqsubsetneq W$, a contradiction. The same argument excludes that $kW\sqsubsetneq W$.
  • If $kW\bot W$, since $\mathcal{C}W$ is unbounded, the partial realisation axiom (10) can be used to find an element $g\in E$ such that $\mathrm{d}_{W}(x_{0},gx_{0})\leq\delta$, while $\mathrm{d}_{kW}(x_{0},gx_{0})\geq M+\delta+1$. This would contradict the fact that $|\mathrm{d}_{W}(x_{0},gx_{0})-\mathrm{d}_{kW}(x_{0},gx_{0})|\leq M$.
  • Finally, suppose $W\pitchfork kW$. Since $\mathcal{C}W$ is unbounded, we can find elements $g,h\in E$ such that
    $\min\{\mathrm{d}_{W}(\rho^{kW}_{W},hx_{0}),\mathrm{d}_{W}(\rho^{kW}_{W},ghx_{0}),\mathrm{d}_{W}(hx_{0},ghx_{0})\}\geq M+3\delta+1.$
    However the consistency axiom (7) would give that
    $\mathrm{d}_{kW}(hx_{0},ghx_{0})\leq\mathrm{d}_{kW}(hx_{0},\rho^{W}_{kW})+\mathrm{d}_{kW}(\rho^{W}_{kW},ghx_{0})+\mathrm{diam}\rho^{W}_{kW}\leq 3\delta,$
    again contradicting that $|\mathrm{d}_{W}(hx_{0},ghx_{0})-\mathrm{d}_{kW}(hx_{0},ghx_{0})|\leq M$.
(2) Let $\mathcal{U}=\bigcup_{i-1}^{n}\operatorname{Big}\left(z_{i}\right)$ and $\widetilde{\mathfrak{S}}=\mathfrak{S}-\mathcal{U}$. By the arguments from [ANS:UEG, Section 3], each $U\in\mathcal{U}$ is orthogonal to any other unbounded domain; in particular, any two $U,U^{\prime}\in\mathcal{U}$ are orthogonal. Notice that $E$ permutes each $\operatorname{Big}\left(z_{i}\right)$, since for every $h\in E$ and $x\in\mathcal{X}$ we have
$\mathrm{diam}\pi_{hU}(\langle z_{i}\rangle\cdot x)=\mathrm{diam}\pi_{U}(h^{-1}\langle z_{i}\rangle\cdot x)=\mathrm{diam}\pi_{U}(\langle z_{i}\rangle h^{-1}\cdot x)=\infty.$
Thus there exists a finite-index, normal subgroup $E^{\prime}\leq E$ fixing every $U\in\mathcal{U}$.
We first prove that a specific $K$-orbit in $\mathcal{X}$ has uniformly bounded projections to all $W\in\widetilde{\mathfrak{S}}$. Complete $\mathcal{U}$ to a maximal collection $\mathcal{V}$ of pairwise orthogonal, unbounded domains, each of which does not contain other unbounded domains. Moreover, let $x_{0}\in\mathcal{X}$ be any point given by the partial realisation axiom (10) applied to the above collection, with any choice of coordinates $p_{i}\in\mathcal{C}V_{i}$. Every other unbounded domain $W\in\widetilde{\mathfrak{S}}$ must contain, or be transverse to, some $V\in\mathcal{V}$, so $\pi_{W}(x_{0})$ is uniformly close to $\rho^{V}_{W}$. But $K$ fixes both $V$ and $W$ by (1), so every $K$-translate of $x_{0}$ projects uniformly close to $\rho^{V}_{W}$. This shows that $\mathrm{diam}\pi_{W}(Z\cdot x_{0})$ is uniformly bounded whenever $W\in\widetilde{\mathfrak{S}}-\mathcal{V}$. Furthermore, for every $V\in\mathcal{V}-\mathcal{U}$, $\langle z_{1}\rangle\cdot x_{0}$ has bounded projection to $\mathcal{C}V$; then $\langle z_{1},z_{2}\rangle\cdot x_{0}=\langle z_{2}\rangle\cdot(\langle z_{1}\rangle\cdot x_{0})$ has bounded projection to $\mathcal{C}V$ as well, and analogously we get that $\mathrm{diam}_{V}K\cdot x_{0}<\infty$. Since there are finitely many $V_{i}$, we just proved that $\sup_{U\in\widetilde{\mathfrak{S}}}\mathrm{diam}\pi_{U}(Z\cdot x_{0})$ is bounded by some constant $C^{\prime}$.
Now notice that
$\sup_{g\in E^{\prime},\,U\in\widetilde{\mathfrak{S}}}\mathrm{diam}\pi_{U}(Z\cdot gx_{0})$ $=\sup_{g\in E^{\prime},\,U\in\widetilde{\mathfrak{S}}}\mathrm{diam}\pi_{g^{-1}U}(Z\cdot x_{0})$
$=\sup_{U\in\widetilde{\mathfrak{S}}}\mathrm{diam}\pi_{U}(Z\cdot x_{0})\leq C^{\prime},$
where we used that every $g\in E^{\prime}$ commutes with $Z$ and preserves $\widetilde{\mathfrak{S}}$. Since $E^{\prime}$ acts coboundedly on $\mathcal{X}$, and since projections are uniformly coarsely Lipschitz, the above bound implies Item (2). ∎
3.3. Bounded central extensions and hierarchical hyperbolicity
We now move to the cohomological characterisation of when a central extension of a HHG, with finitely generated kernel, is a HHG.
Theorem 3.11.
Let $G$ be a HHG, and let $1\to K\to E\to G\to 1$ be a central extension with finitely generated kernel. Then $E$ is a HHG if and only if the extension is bounded.
Remark 3.12.
One direction of Theorem 3.11 extends [uniform_undistortion, Theorem 5.14], which states that a bounded $\mathbb{Z}$-central extension of a HHG is a HHG. In turn, the argument there is a refinement of [HRSS_3manifold, Corollary 4.3], which dealt with the case with hyperbolic base and only relied on the fact that every $\mathbb{Z}$-central extension of a hyperbolic group is bounded [neumannreeves]. As we have seen in Example 2.13, this is no longer automatically true for general HHG. Even assuming acylindrical hyperbolicity of the base is not enough: indeed $H^{2}(\mathbb{Z}^{2}\ast\mathbb{Z};\mathbb{Z})\cong H^{2}(\mathbb{Z}^{2};\mathbb{Z})$, which produces an unbounded class. The argument is the same as Example 2.13, using for instance that the comparison map $H^{2}_{b}(\mathbb{Z}^{2}\ast\mathbb{Z};\mathbb{R})\to H^{2}(\mathbb{Z}^{2}\ast\mathbb{Z};\mathbb{R})$ is trivial [kevin, Example 4.7].
Remark 3.13.
It is natural to ask whether one can also characterise the extension being just HHS rather than HHG. A sufficient, and potentially necessary condition for this is that the extension is quasi-isometrically trivial, which turns out to be equivalent to the Euler class of the extension being weakly bounded, in the sense of [neumannreeves]. For many groups (for instance right-angled Artin groups), weakly bounded classes are bounded [frigsisto], and this is called property QITB (that is, quasi-isometrically trivial implies bounded). However, there are examples of quasi-isometrically trivial extensions which are not bounded [frigsisto, ascarimilizia]. To summarise, two intriguing questions arise from this:
  1. (1)
    If a central extension of an HHG is an HHS, is the extension quasi-isometrically trivial?
  2. (2)
    Do HHGs satisfy QITB? Equivalently, in view of Theorem 3.11, is a quasi-isometrically trivial extension of a HHG also a HHG?
Proof of Theorem 3.11.
Let $K\cong\mathbb{Z}^{n}\times T$ for some $n\geq 0$ and $T$ finite.
$(\Leftarrow)$ Assume that the extension $E\to G$ is bounded. A finite extension of a HHG is a HHG (via the geometric action on the same space). Moreover, a bounded $\mathbb{Z}$-central extension of a HHG is a HHG, as one sees by combining [ABO, Lemma 4.15] and [uniform_undistortion, Theorem 5.14]. So by Corollary 2.10 and induction, we conclude that $E$ is a HHG.
$(\Rightarrow)$ Now assume that $E$ admits a HHG structure, coming from the action on some HHS $(\mathcal{X},\mathfrak{S})$. Our goal is to produce a quasihomomorphism $\Psi\colon E\to K$ which is the identity on $K$, and then we will conclude by Proposition 2.9.
Fix a basepoint $x\in\mathcal{X}$ and a basis $z_{1},\ldots,z_{n}$ for $K$, which we order in such a way that the first $m$ elements generate a subgroup isomorphic to $\mathbb{Z}^{m}$ and the other elements have finite order. Let $\mathcal{U}=\bigcup_{i=1}^{m}\operatorname{Big}\left(z_{i}\right)$, and let $E^{\prime}$ be the finite-index, normal subgroup of $E$ fixing each $U\in\mathcal{U}$, as in the proof of Lemma 3.10. Notice that, by the same Lemma, $K\leq E^{\prime}$; furthermore, $(\mathcal{X},\mathfrak{S})$ is also a HHG structure for $E^{\prime}$.
By arguing as in [ABO, Lemma 4.20], one can show that $\mathcal{C}U$ is a quasiline for every $U\in\mathcal{U}$. Indeed, $K\leq E^{\prime}$ has unbounded orbits on $\mathcal{C}U$, so $E^{\prime}$ does not act elliptically; moreover, the action is not parabolic as it is cobounded (this is because $\pi_{U}$ is coarsely surjective). By the classification of group actions on hyperbolic spaces (see e.g. [ABO, Theorem 4.2]), there must be a loxodromic element $h\in E^{\prime}$. As $K$ commutes with $h$, it must fix the endpoints $h^{+},h^{-}\in\partial\mathcal{C}U$. In particular, the action of $K$ is not parabolic, so some $z\in K$ is a loxodromic isometry of $\mathcal{C}U$ with endpoints $h^{\pm}$. Again, this means that every element in $E^{\prime}$, which commutes with $z$, must fix both $h^{+}$ and $h^{-}$; hence, as the action is cobounded, $\mathcal{C}U$ must be a quasiline.
The above argument also shows that $E^{\prime}$ acts on $\mathcal{C}U$ coboundedly and without inversions. In particular, we can $E^{\prime}$-equivariantly replace each $\mathcal{C}U$ with a Cayley graph for $E^{\prime}$, so that the projection $E^{\prime}\to\mathcal{C}U$ is the identity at the level of vertices.
Now let $x_{0}\in\mathcal{C}U$ correspond to the identity element of $E$. Let $\phi_{U}\colon E^{\prime}\to\mathbb{R}$ be the Busemann quasimorphism associated to the $E^{\prime}$-action on $\mathcal{C}U$, as in Example 2.7. Since the absolute value of $\phi_{U}(g)$ coarsely coincides with the distance in $\mathcal{C}U$ from the basepoint $x_{0}$, we see that $\phi_{U}$ is a quasi-isometry when seen as a map from $\mathcal{C}U$ to $\mathbb{R}$.
Going back to $K$, Lemma 3.10 tells us $K$-orbits are uniformly bounded in every domain outside $\mathcal{U}$. By the Distance Formula [HHS_II, Theorem 4.5], this means that orbit maps quasi-isometrically embed $K$ inside the product $\prod_{U\in\mathcal{U}}\mathcal{C}U$; in turn, since every $\phi_{U}\colon\mathcal{C}U\to\mathbb{R}$ is a quasi-isometry, we get that the map $\Phi\colon K\to\mathbb{R}^{|\mathcal{U}|}$ mapping $k\in K$ to $(\phi_{U}(k))_{U\in\mathcal{U}}$ is a quasi-isometry. Since every homogeneous quasimorphism on the Abelian group $K$ is a homomorphism [frigerio, Corollary 2.12], the image of $\Phi$ must span an $m$-dimensional subspace. Hence, if we identify $\langle z_{1},\ldots,z_{m}\rangle$ with $\mathbb{Z}^{m}\leq\mathbb{R}^{m}$, there is a linear map $M\colon\mathbb{R}^{|\mathcal{U}|}\to\mathbb{R}^{m}$ such that $\Psi\coloneq M\circ\Phi\colon E^{\prime}\to\mathbb{R}^{m}$ is the identity on $\mathbb{Z}^{m}$. Notice that $\Psi$ is a quasihomomorphism as it is a composition of the quasihomomorphism $\Phi$ and a linear map; however, it is not yet the quasihomomorphism we are looking for, since it is only defined on $E^{\prime}$.
To solve the above problem, let $g_{1},\ldots,g_{r}\in E$ be representatives for the cosets of $E^{\prime}$, and define a quasihomomorphism $\Theta\colon E^{\prime}\to\mathbb{R}^{m}$ by mapping $h\in E^{\prime}$ to
$\Theta(h)=\frac{1}{r}\sum_{i=1}^{r}\Psi(ghg^{-1}).$
Notice that $\Theta$ is still the identity on $\mathbb{Z}^{m}$, since the latter is central in $E$; furthermore, by arguing as in e.g. [short_HHG:I, Lemma 3.4], one can see that $\Theta$ is now almost invariant under conjugation by $E$, so let $D>0$ be such that, for every $e\in E$ and $h\in E^{\prime}$, $|\Theta(h)-\Theta(ehe^{-1})|\leq D$.
Finally, let define $\Xi\colon E\to\mathbb{R}^{m}$ by setting $\Xi(g_{i}h)=\Theta(h)$ for every $h\in E^{\prime}$ and $i=1,\ldots,r$. We now prove that $\Xi$ is a quasihomomorphism. Let
$\mathcal{E}=\{e\in E^{\prime}\mid\exists 1\leq i,j,k\leq r\mbox{ s.t. }g_{i}g_{j}=g_{k}e\}.$
Then, for every $g,g^{\prime}\in\{G_{1},\ldots,g_{r}\}$ and every $h,h^{\prime}\in E^{\prime}$, we have
$ghg^{\prime}h^{\prime}=g^{\prime\prime}e((g^{\prime})^{-1}hg^{\prime})h^{\prime},$
where $g^{\prime\prime}\in\{g_{1},\ldots,g_{r}\}$ and $e\in\mathcal{E}$ are such that $gg^{\prime}=g^{\prime\prime}e$. Hence, if the defect of $\Theta$ is $C>0$, we have that
$\left|\Xi(gh)+\Xi(g^{\prime}h^{\prime})-\Xi(ghg^{\prime}h^{\prime})\right|$
$=$ $\left|\Theta(h)+\Theta(h^{\prime})-\Theta(e((g^{\prime})^{-1}hg^{\prime})h^{\prime})\right|$
$\leq$ $\left|\Theta(h)+\Theta(h^{\prime})-\Theta(e)-\Theta((g^{\prime})^{-1}hg^{\prime})-\Theta(h^{\prime})\right|+2C$
$\leq$ $\left|\Theta(h)+\Theta(h^{\prime})-\Theta(e)-\Theta(h)-\Theta(h^{\prime})\right|+2C+D$
$\leq$ $|\Theta(e)|+4C+D$
$\leq$ $\max_{e\in\mathcal{E}}|\Theta(e)|+4C+D.$
Up to taking the integer part, we can assume that $\Xi$ takes values in $\mathbb{Z}^{m}$, and up to a further bounded modification we can assume that $\Xi|_{K}$ is the identity. As a bounded modification of a quasihomomorphism with abelian target is still a quasihomomorphism, $\Xi$ satisfies the requirements of Proposition 2.9, and the proof is complete. ∎
Next, we prove a statement in the opposite direction:
Theorem 3.14.
Let $1\to K\to E\to G\to 1$ be a central extension with finitely generated kernel. If $E$ is a HHG, then the extension is bounded. Moreover, if $E$ has a HHG structure with clean containers, there exists a finite-index subgroup $E^{\prime}\leq E$ containing $K$ such that $E^{\prime}/K$ is a HHG.
Proof.
The proof of $(\Rightarrow)$ of Theorem 3.11 did not use that $G$ is a HHG, so it shows that if $E$ is HHG, then the extension is bounded. We now focus on proving the “moreover” part, whose proof we split into several steps.
Step 1: choice of the finite-index subgroup. Consider a HHG structure $(\mathcal{X},\mathfrak{S})$ for $E$ with clean containers. Fix a basis $\{z_{1},\ldots,z_{n}\}$ for $K$, and let
$\mathcal{U}=\bigcup_{i=1}^{n}\operatorname{Big}\left(z_{i}\right).$
As in the proof of Theorem 3.11, $E$ acts on each $\operatorname{Big}\left(z_{i}\right)$, so let $E^{\prime}$ be the finite-index subgroup of $E$ that fixes each $U\in\mathcal{U}$; notice that $K\leq E^{\prime}$ by Lemma 3.10. Since $E^{\prime}$ has finite index, $(\mathcal{X},\mathfrak{S})$ is a HHG structure for $E^{\prime}$, so we can identify $\mathcal{X}$ with a Cayley graph for $E^{\prime}$ with respect to a finite generating set. For later purposes, let $\delta$ be a HHG constant for the structure.
Step 2: removing unnecessary bounded domains. In this step we restrict the domain set of the HHG structure. Firstly, let
$\mathfrak{S}^{\prime}=\{W\in\mathfrak{S}\mid\exists V\sqsubseteq W\mbox{ s.t. }\mathcal{C}V\mbox{ unbounded}\}.$
We claim that $(E^{\prime},\mathfrak{S}^{\prime})$ is again a HHG structure with clean containers, with the same domains and coordinate spaces. We point out that $\mathfrak{S}^{\prime}$ is $E^{\prime}$-invariant and contains all unbounded domains; therefore, by inspection of Definition 3.1, removing some bounded domains can only affect the existence of containers and the validity of the large link axiom, so we must check that both still hold in $\mathfrak{S}^{\prime}$.
  • Clean containers: Let $V,W\in\mathfrak{S}^{\prime}$ be such that $V\sqsubsetneq W$, and there exists $U\in\mathfrak{S}^{\prime}$ which is nested in $W$ and orthogonal to $V$. By the containers axiom in $\mathfrak{S}$, there exists $T\in\mathfrak{S}$ such that $T\sqsubsetneq W$, $T\bot V$, and $T$ contains all $U$ as above. Since $U\in\mathfrak{S}^{\prime}$, it contains an unbounded domain, and in turn this means that $T\in\mathfrak{S}^{\prime}$. Hence $T$ is the container for $V$ inside $W$ in $\mathfrak{S}^{\prime}$.
  • Large links: By Remark 3.7, it is enough to check the passing up axiom (9’). In turn, since $\mathfrak{S}^{\prime}\subseteq\mathfrak{S}$ has the bounded domain dichotomy, by Remark 3.8 it is enough to check the passing up axiom for unbounded domains, which already held in $\mathfrak{S}$.
[…]
Interlude: Reduction to cyclic kernel. In view of the above discussion, we shall say that a collection $\{h_{1},\ldots,h_{n}\}$ of elements of a HHG $(H,\mathfrak{S})$ is clean if for every $W\in\mathfrak{S}$, $i=1,\ldots,n$, and $U\in\operatorname{Big}\left(h_{1}\right)$, either $U\sqsubseteq W$ or $U\bot W$.
Now let $z=z_{1}$ and $Z=\langle z\rangle$. Our goal is to prove that $E^{\prime}/Z$ admits a HHG structure $\overline{\mathfrak{S}}$ such that the image of $\{z_{2},\ldots,z_{n}\}$ is clean in $(E^{\prime}/Z,\overline{\mathfrak{S}})$. Therefore, if in the following arguments we are careful to never use that $\mathfrak{S}^{\prime\prime}$ has clean containers, but only that $\{z_{1},\ldots,z_{n}\}$ is clean in $(E^{\prime},\mathfrak{S}^{\prime\prime})$, then we will be able to conclude by induction that $E^{\prime}/K$ is a HHG, thus proving Theorem 3.14
Step 3: reduction to a single big domain. In this step we replace some coordinate spaces in $\mathfrak{S}^{\prime\prime}$ to get a new HHG structure where $z$ has at most one big domain (this step is unnecessary if $z$ has finite order). Let $\operatorname{Big}\left(z\right)=\{U_{1},\ldots,U_{j}\}$, and for every $i=1,\ldots,j$ let $\phi_{i}\coloneq E^{\prime}\to\mathbb{R}$ be the Busemann quasimorphism for the action on $\mathcal{C}U_{i}$. We first claim that these quasimorphisms are linearly independent. Indeed, if, say, $\phi_{1}$ was a linear combination of the others, then there would be some function $f\colon\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ such that $\mathrm{d}_{U_{1}}(g,1)\leq f(\mathrm{d}_{U_{2}}(g,1),\ldots,\mathrm{d}_{U_{j}}(g,1))$ for every $g\in E^{\prime}$ (here we used that, as argued in Example 2.7, the absolute value of each quasimorphism coarsely coincides with the distance in the corresponding quasiline). However the elements of $\operatorname{Big}\left(z\right)$ are pairwise orthogonal, so the partial realisation axiom (10) implies that $E^{\prime}$ acts coboundedly on the product $\prod_{i=1}^{j}\mathcal{C}U_{i}$. In particular, one can find a sequence of $g_{n}\in E^{\prime}$ for which $\mathrm{d}_{U_{1}}(g_{n},1)\to\infty$ while the other projections are uniformly bounded, contradicting the existence of $f$.
Next, set $\psi_{1}=\phi_{1}$, and for every $2\leq i\leq j$ let $\psi_{i}=\phi_{i}-\frac{\phi_{i}(z)}{\phi_{1}(z)}\phi_{1}$, which is non-trivial by linear independence. For every $1\leq i\leq j$ let $L_{i}$ be a Cayley graph for $E^{\prime}$ such that $\psi_{i}\colon L_{i}\to\mathbb{R}$ is a quasi-isometry, as constructed in [ABO, Lemma 4.15]. Notice that $\psi_{i}(z)=0$ whenever $i\geq 2$, so $Z$ now acts elliptically on all quasilines excluding $L_{1}$.
Now, replacing $\mathcal{C}U_{1}$ by $L_{1}$ gives a new HHG structure for $E^{\prime}$, since the two spaces are $E^{\prime}$-equivariantly quasi-isometric. We then replace each $\mathcal{C}U_{i}$ by the corresponding $L_{i}$, with projection given by the identity map $\mathcal{X}\to L_{i}$ as both are Cayley graphs for $E^{\prime}$. Notice that it is not necessary to define new relative projections, as no domain is transverse to, or nested in, any $U_{i}$.
We claim that the above replacement procedure gives a HHG structure for $E^{\prime}$, where $z$ will then have a single big domain. All requirements only involving the relations on $\mathfrak{S}^{\prime\prime}$, or the projections between domains in $\mathfrak{S}^{\prime\prime}$, still hold by construction, so we focus on the axioms which impose restrictions on the $L_{i}$.
  • Projections (1). The projection $\mathcal{X}\to L_{i}$ is surjective and Lipschitz, as $\mathcal{X}$ is a Cayley graph for $E^{\prime}$ with respect to a finite generating set.
  • Consistency (7). Since $\{z\}$ is clean in $\mathfrak{S}^{\prime}$, no element of $\mathfrak{S}^{\prime\prime}$ is either transverse to, or nested in, any $U\in\operatorname{Big}\left(z\right)$. Hence, both consistency inequalities only involve domains from $\mathfrak{S}^{\prime\prime}-\operatorname{Big}\left(z\right)$, and therefore already held in the original structure.
  • Bounded geodesic image (8). Since every $U_{i}$ is orthogonal to every other unbounded domain, there are no pairs of nested unbounded domains involving $U_{i}$; all other instances of the axiom already appeared in the original structure.
  • Large links (9). As we only replaced finitely many coordinate spaces, the axiom still holds (possibly after replacing $\delta$ by a bigger constant).
  • Partial realisation (10). Let $\mathcal{V}=\{(V_{i},q_{i}\in\mathcal{C}V_{i})\}_{i=1,\ldots,m}$ be a collection of pairwise orthogonal domains, each with a point in the associated coordinate space. If no domain in $\mathcal{V}$ belongs to $\operatorname{Big}\left(z\right)=\emptyset$ then the original partial realisation axiom already provided a realisation point, so suppose otherwise. If $U_{1}$ appears in $\mathcal{V}$ then let $g_{1}$ be the coordinate $q_{1}\in L_{1}$ that we have to realise; otherwise let $g_{1}=1\in E^{\prime}$. Whenever $U_{i}\in\mathcal{V}$ let $q_{i}\in L_{i}$ the coordinate to realise, choose $g_{i}\in E^{\prime}$ such that $\phi_{i}(g_{i})=\psi_{i}(q_{i})+\psi_{1}(g_{1})$, and let $p_{i}=\pi_{\mathcal{C}U_{i}}(g_{i})$. By the partial realisation axiom for the original structure, there exists some $x\in\mathcal{X}$ realising the collection
    $\{(U_{i},p_{i}\in\mathcal{C}U_{i})\}_{U_{i}\in(\operatorname{Big}\left(z\right)-\{U_{1}\})\cap\mathcal{V}}\cup\{(V_{i},q_{i}\in\mathcal{C}V_{i})\}_{V_{i}\in\mathcal{V}-(\operatorname{Big}\left(z\right)-\{U_{1}\})},$
    and we claim that $x$ is uniformly close to $p_{i}$ in each $L_{i}$. Since $x$ is a realisation point, for every $i=1,\ldots,j$ we have that $\mathrm{d}_{\mathcal{C}U_{i}}(g_{i},x)\leq\delta$, so that $\phi_{i}(x)$ and $\phi_{i}(g_{i})$ are uniformly close. In turn, this means that $\psi_{i}(x)$ is uniformly close to $\psi_{i}(p_{i})$ for all $i$, and therefore $\mathrm{d}_{L_{i}}(p_{i},x)$ is uniformly bounded, as required.
  • Uniqueness (11) Suppose $x,y\in\mathcal{X}$ have uniformly close projections to all domains in the new structure. If we show that their projections to each $\mathcal{C}U_{i}$ are also uniformly bounded then $x$ and $y$ are uniformly close, by the uniqueness axiom in the original structure. Up to the group action we can assume that $y=1$, so that $\mathrm{d}_{L_{i}}(1,x)$ coarsely coincides with the absolute value of $\psi_{i}(x)$. Since all distances in the $L_{i}$ are uniformly bounded, so are the absolute values of $\psi_{i}$, hence of $\phi_{i}$, and again this means that $\mathrm{d}_{\mathcal{C}U_{i}}(1,x)$ is uniformly bounded for all $i$.
  • HHG structure. All projections (old and new) are $E^{\prime}$-equivariant, so we built a hierarchically hyperbolic group structure.
Before proceeding, we point out that the collection $\{z_{2},\ldots,z_{n}\}$ is still clean in the new structure. This is because we did not change the projections to any domain outside $\mathcal{U}$; hence, for every $i=2,\ldots,n$, $\operatorname{Big}\left(z_{i}\right)$ might consist of different domains but it must still be contained inside $\mathcal{U}$.
Step 4: HHG structure of the quotient. We are finally ready to prove that $E^{\prime}/Z$ is a HHG, with the structure that we now describe. If $z$ has infinite order let $U_{z}$ be the unique big domain for $z$ in the new structure, and let $\widetilde{\mathfrak{S}}=\mathfrak{S}^{\prime\prime}-\{U_{z}\}$; otherwise let $\widetilde{\mathfrak{S}}=\mathfrak{S}^{\prime\prime}$. Let $\mathcal{Y}=\mathcal{X}/Z$ and $\overline{\mathfrak{S}}=\widetilde{\mathfrak{S}}/Z$. Given a domain $W\in\widetilde{\mathfrak{S}}$, let $[W]$ be its image in $\overline{\mathfrak{S}}$. Two domains $U],[V]\in\overline{\mathfrak{S}}$ are orthogonal (resp. nested) if they admit orthogonal (resp. nested) representatives $U\in[U]$ and $V\in[V]$. Set
$\mathcal{C}[W]=\left(\bigcup_{W\in[W]}\mathcal{C}W\right)/K,$
and for any representative let $q_{W}\colon\mathcal{C}W\to\mathcal{C}[W]$ be the quotient projection. This map is $1$-Lipschitz; moreover, it is a uniform quality quasi-isometry, because $K$-orbits are uniformly bounded in $\mathcal{C}W$ by Lemma 3.10. In particular, the spaces $\mathcal{C}W$ are uniformly hyperbolic.
Given $[W]$, if $\mathcal{C}W$ is bounded then so is $\mathcal{C}[W]$, and projections can be defined arbitrarily. Otherwise there is a single representative $W\in[W]$, and given $x]\in\mathcal{Y}$ let
$\pi_{[W]}([x])=q_{W}\left(\bigcup_{x\in[x]}\pi_{W}(x)\right).$
Similarly, for every $V]\in\overline{\mathfrak{S}}$ which is properly nested in, or transverse to, $[W]$, set
$\rho^{[V]}_{[W]}=q_{W}\left(\bigcup_{V\in[V]}\rho^{V}_{W}\right).$
We now check that $(\mathcal{Y},\overline{\mathfrak{S}})$ satisfies the axioms of an almost HHG structure.
  • Projections (1). For every $W]\in\overline{\mathfrak{S}}$, $\pi_{[W]}$ is a $\delta$-coarsely onto, $\delta$-coarse map, as the quotient projection $q_{W}\colon\mathcal{C}W\to\mathcal{C}[W]$ is $1$-Lipschitz and surjective. Moreover, given $x],[y]\in\mathcal{Y}$ let $x\in[x]$ and $y\in[y]$ realise the distance between $[x]$ and $[y]$, so that
    $\mathrm{d}_{[W]}([x],[y])\leq\mathrm{d}_{W}(x,y)\leq\delta\mathrm{d}_{\mathcal{X}}(x,y)+\delta=\delta\mathrm{d}_{\mathcal{Y}}([x],[y])+\delta.$
  • Nesting (2). By construction, the unique maximal element of $\overline{\mathfrak{S}}$ is $[S]$. Moreover, whenever $V]\sqsubsetneq[W]$ and $\mathcal{C}W$ is unbounded, we have that $\mathrm{diam}\rho^{[V]}_{[W]}\leq\mathrm{diam}\bigcup_{k\in Z}\rho^{kV}_{W}$, and the latter is uniformly bounded since $Z$-orbits in $\mathcal{C}W$ are uniformly bounded.
  • Finite complexity (3). By how nesting is defined, every chain $U_{1}]\sqsubsetneq\ldots\sqsubsetneq[U_{k}]$ in $\overline{\mathfrak{S}}$ lifts to a chain $U_{1}\sqsubsetneq\ldots\sqsubsetneq U_{k}$ in $\widetilde{\mathfrak{S}}$, whose length is uniformly bounded.
[…]
    Moving to the second requirement, let $U],[V],[W]\in\overline{\mathfrak{S}}$ be such that $U]\sqsubsetneq[V]$ and $V]\bot[W]$. As above, we can find representatives $U\sqsubsetneq V$ and $V\bot W$, so we get that $U\bot W$ and therefore $U]\bot[W]$.
  • Containers (5). Let $V]\sqsubsetneq[W]$, and let $V^{\prime}]\sqsubseteq[V]$ be such that $\mathcal{C}V^{\prime}$ is unbounded. Let $\mathcal{Q}=\{[Q]\sqsubsetneq[W]\mid[Q]\bot[V]\}$, and notice that every element of $\mathcal{Q}$ is orthogonal to $V^{\prime}]$ as well. Fix a representative $W\in[W]$, and for every $Q]\in\mathcal{Q}$ let $Q\in[Q]$ be nested in $W$. Since $V^{\prime}$ is the unique representative of $V^{\prime}]$, it must be nested in $W$ and orthogonal to every $Q$. Thus the container axiom for $\widetilde{\mathfrak{S}}$ produces a domain $T\sqsubsetneq W$ which contains every $Q$. Hence $T]\sqsubsetneq[W]$ and contain all $Q]\in\mathcal{Q}$, so it is a container for $[V]$ inside $[W]$.
  • Transversality (6). Uniform boundedness of projection points follows as for the nesting axioms.
[…]
  • Partial realisation (10). Let $V_{1}],\ldots,[V_{k}]\in\overline{\mathfrak{S}}$ be pairwise orthogonal, and for every $i$ let $p_{i}\in\mathcal{C}[V_{i}]$. For every $i$ let $r_{i}\in q_{V_{i}}^{-1}(p_{i})$, and let $x\in\mathcal{X}$ realise the collection $\{r_{i}\}_{i=1,\ldots,k}$. It is easy to see that $[x]$ realises $\{p_{i}\}_{i=1,\ldots,k}$ in $\mathcal{Y}$.
[…]
  • HHG structure. Since $E^{\prime}$ acted metrically properly and coboundedly on $\mathcal{X}$, then so does $E^{\prime}/Z$ on $\mathcal{Y}$. Moreover, the cofinite $E^{\prime}$-action on $\widetilde{\mathfrak{S}}$ induces a cofinite $E^{\prime}/Z$-action on $\overline{\mathfrak{S}}$, so $(\mathcal{Y},\overline{\mathfrak{S}})$ is a HHG structure for $E^{\prime}/Z$.
Step 5: conclusion As argued in the Interlude above, to conclude the proof of Theorem 3.14 we are left to show that the collection $\{[z_{2}],\ldots,[z_{n}]\}\subset E^{\prime}/K$ is clean in $\overline{\mathfrak{S}}$. By how orthogonality and nesting are defined in $\overline{\mathfrak{S}}$, it is in turn enough to prove that $\operatorname{Big}\left([z_{i}]\right)\subseteq[\mathcal{U}]$, where $\mathcal{U}]=\{[U]\}_{U\in\mathcal{U}-\{U_{z}\}}$. Indeed, if $W]\in\overline{\mathfrak{S}}$ has unbounded coordinate space, and $\mathrm{diam}_{[W]}(\langle[z_{i}]\rangle\cdot[x_{0}])=\infty$ for some $x_{0}\in\mathcal{X}$, then $\mathrm{diam}_{W}(\langle z_{i}\rangle\cdot x_{0})=\infty$ as well, since the projection $q_{W}\colon\mathcal{C}W\to\mathcal{C}[W]$ is a quasi-isometry. Hence $W\in\operatorname{Big}\left(z_{i}\right)-\{U_{z}\}$, so $W]\in[\mathcal{U}]$, as required. ∎
Remark 3.15.
In the statement of Theorem 3.14, one cannot hope that $G$ is genuinely a HHG. Indeed, the $(3,3,3)$ triangle group $G$ is not a HHG [PS_unbounded, Corollary 4.5]; however the direct product $G\times\mathbb{Z}$ is a HHG with clean containers, as it acts geometrically on the standard cubulation of $\mathbb{R}^{3}$. We are thankful to Mark Hagen for explaining to us a proof of the latter fact, which we now present. Let $o=(0,0,0)$ and $p=(1,1,1)$. Let $H_{1}$ be the plane $\{x=y\}$, and similarly define $H_{2}=\{y=z\}$ and $H_{3}=\{x=z\}$. Let $T_{1},T_{2},T_{3}$ be the reflections across $H_{1},H_{2},H_{3}$, which preserve the standard cubulation of $\mathbb{R}^{3}$. Now let $F=\{x+y+z=0\}$, which is the plane orthogonal to $op$ passing through $o$, and let L(x,y,z)=(x+1,y+1,z2)L(x,y,z)=(x+1,y+1,z-2), which is an integer translation along the line in which $H_{1}$ intersects $F$. Notice that $T_{1},T_{2},T_{3}$ preserve $F$ as they all fix $op$, and moreover so does $L$. Then let $G=\langle\alpha,\beta,\gamma\rangle$, where
  • $\alpha=T_{1}\colon(x,y,z)\mapsto(y,x,z)$,
  • $\beta=T_{2}\colon(x,y,z)\mapsto(x,z,y)$,
  • $\gamma=LT_{3}L^{-1}\colon(x,y,z)\mapsto(z-3,y,x+3)$.
$G$ is a quotient of the $(3,3,3)$ triangle group $H$, since the above elements are all reflections and the product of any two of them has order $3$ (geometrically, this is because the dihedral angle between any two of the planes is $\pi/3$). Also $\alpha\circ\beta\circ\gamma$ is given by $(x,y,z)\mapsto(x+3,z-3,y)$, which has infinite order (as one sees by looking at the first coordinate). However, $H$ is just-infinite, meaning that its only infinite quotient is $H$ itself, so $G\cong H$. Indeed, this follows from [McCarthy_just_infinite, Proposition 9], together with the fact that the point group is $S_{3}$, realised as an irreducible subgroup of $\mathrm{GL}_{2}(\mathbb{Z})$.
Finally, let $S(x,y,z)=(x+1,y+1,z+1)$, which commutes with $L$ and with every $T_{i}$ as $op\subset H_{1}\cap H_{2}\cap H_{3}$. Moreover $\langle S\rangle\cap G$ is trivial, since $G$ fixes $F$. Hence $E=\langle S,G\rangle\cong\langle S\rangle\times G$ is a direct product. Notice that $E$ preserves the standard cubulation of $\mathbb{R}^{3}$, as its generators do, and the action is proper and cocompact (as one can see by considering the plane $F$ and its translates by powers of $S$). By e.g. [HHS_I, Remark 13.2], $E$ is a HHG, whose domain set consists of:
  • the top element $S$;
  • three elements for the coordinate planes, which are pairwise transverse;
  • three elements for the coordinate lines, each of which is nested in the two planes it belongs and orthogonal to the third one.
In particular this structure has clean containers, as required.
4. Quotients of mapping class groups
4.1. The general problem
We now discuss certain quotient extensions that it would be interesting to understand. The context is quotients of mapping class groups, and in particular the conjecture of the second and third author [MS_rigidity_MCG, Question 3], which we now state in a simplified form.
Let $\mathcal{MCG}(S)$ be the mapping class group of a finite-type surface $S$, and let $\mathcal{B}=\{\phi_{1},\dots,\phi_{k}\}$ be a collection of elements. The conjecture predicts that there exists $M\in\mathbb{N}_{>0}$ such that $\mathcal{MCG}(S)/\mathcal{N}$ is hierarchically hyperbolic, where
$\mathcal{N}=\langle\hskip-1.99168pt\langle\phi_{1}^{M},\ldots,\phi_{k}^{M}\rangle\hskip-1.99168pt\rangle.$
There are various cases known in the literature. For instance, [BHMS, Theorem 7.1] provides an affirmative answer if $\mathcal{B}$ consists of conjugacy representatives of all Dehn twists; [short_HHG:II] almost completely settles the conjecture for the five-punctured sphere; and [random_quot] proves hierarchical hyperbolicity of quotients by random walks, which should be thought as the “typical” quotients (see also [hhs_asdim] for similar classes of quotients). The simplest unknown case is where $\mathcal{B}$ consists of a single Dehn twist, and even in that case one already encounters the problem of showing that a certain quotient central extension is bounded.
More precisely, if $p\colon\mathcal{MCG}(S)\to E$ is the quotient under consideration, let $H\leq E$ be the image under $p$ of the stabiliser of a curve $\gamma$, with the property that no power of the Dehn twist around $\gamma$ is conjugate into an element of $\mathcal{B}$. In any reasonable HHG structure, $H$ should be itself hierarchically hyperbolic, as it would correspond to a standard product region (see e.g. [HHS_II, Definition 5.15]). One of the simplest scenarios is where $\gamma$ is non-separating, so that its stabiliser is the mapping class group of $S-\gamma$. Hence, a simplified, yet significant version of the problem is the following.
Question 4.1.
Let $U$ be a finite-type surface with one boundary component $\gamma$, let $\widehat{U}$ be the surface obtained by gluing a once-punctured disk to $\gamma$ with puncture $p$, and let $\mathcal{MCG}(\widehat{U},p)$ be the subgroup of the mapping class group of $\widehat{U}$ fixing $p$. There is a short exact sequence
$1\to\langle\tau_{\gamma}\rangle\to\mathcal{MCG}(U)\to\mathcal{MCG}(\widehat{U},p)\to 1.$
Let $\mathcal{N}_{\gamma}\leq\mathcal{MCG}(U)$ be the normal subgroup generated by all $M$-th powers of Dehn twists along curves which are supported on the interior of $U$. Is $\mathcal{MCG}(U)/\mathcal{N}_{\gamma}$ hierarchically hyperbolic, for a suitable choice of $M$?
It can be deduced from [dahmani:rotating] that $\mathcal{N}_{\gamma}$ intersects the kernel trivially, so the quotient is itself a central extension
$1\to\langle\tau_{\gamma}\rangle\to\mathcal{MCG}(U)/\mathcal{N}_{\gamma}\to\mathcal{MCG}(\widehat{U},p)/\pi(\mathcal{N}_{\gamma})\to 1.$
Furthermore, the aforementioned [BHMS, Theorem 7.1] yields that the base $\mathcal{MCG}(\widehat{U},p)/\pi(\mathcal{N}_{\gamma})$ is itself hierarchically hyperbolic, for a suitable choice of $M$; hence Question 4.1 is equivalent to the boundedness of the second central extension, by Theorem 3.11.
4.2. Quotients by pseudo-Anosovs, low complexity
Another interesting (and possibly easier) version of Question 4.1 is to ask about quotients by powers of pseudo-Anosovs rather than Dehn twists.
Question 4.2.
Let $U$ be a finite-type surface with one boundary component $\gamma$, and let $\widehat{U}$ be the surface obtained by gluing a once-punctured disk to $\gamma$ with puncture $p$. Let $h$ be a pseudo-Anosov mapping class, and let $\mathcal{N}_{h}$ be the normal closure of a sufficiently high power of $h$. Is the central extension
$1\to\langle\tau_{\gamma}\rangle\to\mathcal{MCG}(U)/\mathcal{N}_{h}\to\mathcal{MCG}(\widehat{U},p)/\pi(\mathcal{N}_{h})\to 1$
hierarchically hyperbolic? Equivalently, since the base is hierarchically hyperbolic by [hhs_asdim, Theorem 6.2], is the extension bounded?
In this section we answer the question in the case where $U$ is a 4-punctured disk, whose mapping class group is the braid group on 4 strands, and therefore $\widehat{U}$ is a 5-punctured sphere $S_{5}$. As we will see, it is crucial for the proof that modding out powers of Dehn twists in the mapping class group of the five-punctured sphere yields a hyperbolic group. Later, we also discuss a different strategy to approach the general case.
Theorem 4.3.
Let $h$ be a pseudo-Anosov element of $B_{4}$, and let $\hat{h}$ be its image in $\mathcal{MCG}(S_{5},p)$. Then there exists $M_{0}\in\mathbb{N}_{>0}$ such that for all multiples $M$ of $M_{0}$ the following holds. Let $\mathcal{N}_{h}\unlhd B_{4}$ (resp. $\mathcal{N}\unlhd\mathcal{MCG}(S_{5},p)$) be the normal subgroup generated by $h^{M}$ (resp. $\hat{h}^{M}$), with quotient map $q$ (resp. $\widehat{q}$). Then there is a commutative diagram
[…]
where the bottom row is a bounded central extension of a HHG. In particular, $B_{4}/\mathcal{N}_{h}$ is hierarchically hyperbolic.
One can understand subgroups generated by powers of pseudo-Anosovs using the technology of rotating families; this is done in [DGO]. To prove the theorem, we will need an improvement of this technology from [dahmani:rotating]. We do not recall all the relevant definitions here, but we discuss the main points, giving the motivating example from [dahmani:rotating] along the way. A composite projection system is a set $\mathbb{Y}_{\ast}$ (for instance, the set of curves on a surface) together with additional data. One piece of data is, for each $X\in\mathbb{Y}_{\ast}$, a subset $Act(X)$ of $\mathbb{Y}_{\ast}$ (for instance, $Act(X)$ is the set of curves that intersect $X$). A composite rotating family on a composite projection system is given by groups $G_{X}$ of automorphisms of $\mathbb{Y}_{\ast}$ fixing $X$, for all $X\in\mathbb{Y}_{\ast}$, satisfying certain conditions (for instance, $G_{X}$ could be the subgroup generated by the Dehn twist around $X$).
We need the following technical proposition to make sure that the central extensions we will consider are still $\mathbb{Z}$-central extensions.
Proposition 4.4.
Let $G$ be a group acting by automorphisms on a composite projection system $\mathbb{Y}_{\ast}$, and let $Z$ be the kernel of the action. Suppose that we have subgroups $\{G_{X}\}_{X\in\mathbb{Y}_{\ast}}$ of $G$ such that the following hold.
  • $Z\cap G_{X}=\{1\}$.
  • $G_{gX}=gG_{X}g^{-1}$.
  • $G_{X}$ commutes with $G_{Y}$ if $X\notin Act(Y)$.
  • denoting $\hat{G}_{X}$ the image of $G_{X}$ in $G/Z$, we have that $\{\hat{G}_{X}\}$ forms a composite rotating family on $\mathbb{Y}_{\ast}$.
Then the quotient map $p\colon G\to G/Z$ restricts to an isomorphism
$N=\langle\hskip-1.99168pt\langle\{G_{X}\}\rangle\hskip-1.99168pt\rangle_{G}\to\langle\hskip-1.99168pt\langle\{\hat{G}_{X}\}\rangle\hskip-1.99168pt\rangle_{G/Z}=\hat{N}$.
Proof.
Clearly we have $p(N)=\hat{N}$, so we can restrict $p$ to $N$ to obtain a surjective homomorphism $\bar{p}\colon N\to\hat{N}$. In order to show that this is an isomorphism, we construct an inverse $\iota:\hat{N}\to N$, using the first presentation of $\hat{N}$ provided by [dahmani:rotating, Theorem 2.2], which applies by the fourth bullet. The $\hat{G}_{X}$ are generators for this presentation, and thus have natural maps to the $G_{X}$ by the first bullet. The relations for this presentation are conjugation relations, which are respected in view of the second bullet, and commutation relations, which are respected by the third bullet. Therefore, there is a well-defined homomorphism $N\to\hat{N}$, which when restricted to any $\hat{G}_{X}$ gives an inverse of $\bar{p}|_{G_{X}}$, and therefore $\iota$ is in fact the inverse of $\bar{p}$, as required. ∎
Proof of Theorem 4.3.
Throughout this this proof we always assume that $M$ is a multiple of some suitable $M_{0}\in\mathbb{N}_{>0}$.
First of all, we need $\mathcal{N}_{h}$ to intersect $\langle\tau_{\gamma}\rangle$ trivially, justifying that there is indeed a diagram as in the statement, with the vertical arrow on the left being an isomorphism. This follows from Proposition 4.4, with $G=B_{4}$, $Z=\langle\tau_{\gamma}\rangle$, and $G_{X}$ the conjugates of $\langle h^{M}\rangle$.
In order to show boundedness of the relevant central extension, we now argue that there is a further quotient central extension which is a central extension of a hyperbolic group. More precisely, we argue that there is a diagram
${1}{\mathbb{Z}}$$B_{4}/\mathcal{N}_{h}$$\mathcal{MCG}(S_{5},p)/\mathcal{N}$$1{1}{\mathbb{Z}}$$H^{\hat{H}}$${1,}\scriptstyle{\cong}$
where $\hat{H}$ is hyperbolic. Since the comparison map for hyperbolic groups is surjective [neumannreeves] (see also [mineyev]), by Lemma 2.12 we then get that the central extension from the statement is bounded, as required.
The required quotient $H$ can be taken to be the quotient of $G=B_{4}/\mathcal{N}_{h}$ by the normal subgroup $\mathcal{K}$ generated by suitable powers of all images of Dehn twists of $B_{4}$. We now explain why we can choose such a $\mathcal{K}$ that intersects trivially the image $Z$ of $\langle\tau_{\gamma}\rangle$ in $G$, and such that $G/\mathcal{K}Z$ is hyperbolic. This uses results from [short_HHG:I, short_HHG:II] about short HHG. We do not need the whole definition here; what is important is that these are HHG with a specified collection of cyclic subgroups, and if all of those are finite then the group is hyperbolic, see [short_HHG:I, Lemma 2.14]. All relevant short HHG below also satisfy the assumption called colourability from [DMS:stable, Definition 2.8], which from now on we omit. Modding out suitable powers of generators of any subcollection of the specified cyclic subgroups yields another short HHG by [short_HHG:II, Theorem 4.1], with associated collection of specified subgroups simply the images of the specified subgroups of the original short HHG.
First of all, $\hat{G}=\mathcal{MCG}(S_{5},p)/\mathcal{N}$ is a short HHG. This can be seen combining the following:
  • $\mathcal{MCG}(S_{5},p)$ is a short HHG, see [short_HHG:I, Subsection 2.3.1], with specified subgroups generated by Dehn twists.
  • We can add any cyclic subgroup generated by a pseudo-Anosov to the collection of specified subgroups of the short HHG structure of $\mathcal{MCG}(S_{5},p)$ by [short_HHG:I, Proposition 5.1].
  • The aforementioned [short_HHG:II, Theorem 4.1] allows us to conclude that modding out suitable powers of $h$ yields a short HHG (with the expected structure with specified subgroups coming from images of Dehn twists).
Now, for $\mathcal{K}$ as above we can consider its image $\hat{\mathcal{K}}$ in $\hat{G}$, and this is the kernel associated to a composite rotating family, see [short_HHG:II, Section 3.2]. This allows us to apply Proposition 4.4 and deduce that $\mathcal{K}$ intersects $Z$ trivially, as required. Moreover, $G/{\mathcal{K}}Z=\hat{G}/\hat{\mathcal{K}}$ is a short HHG by [short_HHG:II, Theorem 4.1], and all its specified subgroups are finite, which as mentioned above implies that the group is hyperbolic. This concludes the proof. ∎
Remark 4.5.
Under the assumption that “enough” hyperbolic groups are residually finite, [BHMS] constructs hyperbolic quotients of mapping class groups. One could in principle try to exploit these quotients, or variations, and a similar strategy as in the proof above to solve other cases of Question 4.2, or to solve it completely, but conditionally on the residual finiteness of the relevant hyperbolic groups. The main difficulty is ensuring that kernels to these hyperbolic quotients lift isomorphically to the corresponding central extension, as in Proposition 4.4.
4.3. Quotients by pseudo-Anosov, high complexity?
In the setting of Question 4.2, results from [DGO] allow us to write $\pi(\mathcal{N}_{h})$ as the directed union of groups $N_{i}$, each of which is a free product of finitely many conjugates of the cyclic group generated by the relevant power of $\hat{h}$. Each of the $N_{i}$ is convex-cocompact, and we expect that a variant of Theorem 2.19 allows one to extend any quasimorphism on $N_{i}$ to $\mathcal{MCG}(U)/\pi(\mathcal{N}_{h})$. If this extension could be realised in a way that the defect is uniformly controlled, then we would have a positive answer to the previous question, thanks to the following lemma:
Lemma 4.6.
Let $G$ be a finitely generated group, $N\leq G$ a normal subgroup, and $\chi\colon N\to\mathbb{Z}$ a quasimorphism. Suppose that $N$ can be expressed as a directed union of subgroups $\{N_{i}\}_{i\in I}$, and that for each $i\in I$ there is a quasimorphism $X_{i}\colon G\to\mathbb{Z}$ that extends $\chi|_{N_{i}}$, such that the defect of the $X_{i}$ is uniformly bounded. Then $\chi$ extends to a quasimorphism $X\colon G\to\mathbb{Z}$.
Proof.
By using homogeneous representatives for one direction, and integer parts for the other, as we did before, we can reduce to the following statement. Let $\chi\colon N\to\mathbb{R}$ be a homogeneous quasimorphism. Suppose that $N$ can be expressed as a directed union of subgroup $\{N_{i}\}_{i\in I}$, and that for each $i\in I$ there is a homogeneous quasimorphism $X_{i}\colon G\to\mathbb{R}$ that extends $\chi|_{N_{i}}$, such that the defect of the $X_{i}$ is uniformly bounded. Then $\chi$ extends to a homogeneous quasimorphism $X\colon G\to\mathbb{R}$. Moreover, because $G$ is finitely generated, thus countable, we can assume that the directed set $I$ is just $\mathbb{N}$ with its well order.
Let us make a further reduction: it suffices to show that we can modify each $X_{i}$ to obtain $X_{i}^{\prime}\colon G\to\mathbb{R}$, which still extends $\chi|_{N_{i}}$ and has uniformly bounded defect, and moreover, for every $g\in G$, $\{|X_{i}^{\prime}(g)|\}_{i\in\mathbb{N}}$ is uniformly bounded. Indeed, assuming this, fix a non-principal ultrafilter $\omega$ on $\mathbb{N}$. Then we can set $X(g)\coloneqq\lim_{\omega}X_{i}(g)$, which is a well-defined real number. Moreover, it is a homogeneous quasimorphism, since ultralimits are norm non-increasing and linear.
Now consider the map $\alpha\colon G\to H_{1}(G;\mathbb{Q})$: this factors every homomorphism from $G$ to $\mathbb{Q}$, and its kernel coincides with the set of elements with a power that belongs to $[G,G]$. In particular, if $g\in\ker(\alpha)$, then its stable commutator length $\mathrm{scl}(g)$ is finite, and by Bavard duality [bavard] we have a bound $|X_{i}(g)|\leq 2\mathrm{scl}(g)D(X_{i})$. This shows that for $g\in\ker(\alpha)$, we do not need to change $X_{i}$; note that this concludes the proof in the case in which $G$ has finite abelianisation (e.g. for mapping class groups).
For the general case, let $x_{1},\ldots,x_{n}$ be elements of $G$ that are mapped to a basis of $H_{1}(G;\mathbb{Q})$, chosen in such a way that $x_{1},\ldots,x_{m}\notin N$ while $x_{m+1},\ldots,x_{n}\in N$. Now, for each $j=1,\ldots,m$, let $\lambda_{j}\colon G\to\mathbb{Q}$ be the functional dual to $x_{j}$. For each $i\in\mathbb{N}$, we set
$X_{i}^{\prime}\coloneqq X_{i}-\sum\limits_{j=1}^{m}X_{i}(x_{j})\cdot\lambda_{j}.$
As a result, $X_{i}^{\prime}$ is a homogeneous quasimorphism with the same defect as $X_{i}$, the same restriction on $N$ (in particular it is still an extension of $\chi|_{N_{i}}$) and moreover it vanishes on $x_{j}$ for all $j=1,\ldots,m$. Now for an element $g\in G$, we can write $\alpha(g)=\sum a_{j}\alpha(x_{j})$, for some $a_{j}\in\mathbb{Q}$. Thus, for a high enough power $p$, we can write
$\alpha(g^{p})=\alpha\left(\prod\limits_{j=1}^{n}x_{j}^{p_{j}}\right)$
for some integers $p_{j}\in\mathbb{Z}$. This way we have an expression $g^{p}=xyz$, where $x$ is a product of powers of $x_{1},\ldots,x_{m}$; $y$ belongs to $N$, and $z$ belongs to $\ker(\alpha)$. By the first case treated before, $|X_{i}^{\prime}(z)|$ is uniformly bounded. Moreover, $|X_{i}^{\prime}(y)|$ is uniformly bounded because $y\in N_{i}$ for $i$ large enough, at which point $X_{i}(y)=\chi(y)$. Finally, $|X_{i}^{\prime}(x)|$ is uniformly bounded, since $x$ is a product of $m$ terms, each of which vanish under $X_{i}^{\prime}$. All in all, this shows that $|X_{i}^{\prime}(g^{p})|$ is uniformly bounded, and thus $|X_{i}(g)|=\frac{1}{p}|X_{i}(g^{p})|$ is uniformly bounded, which concludes the proof. ∎
As pointed out in Remark 2.20, in our setting there is a single bounded cocycle that extends $\delta\chi_{i}|_{N_{i}}$. It could be possible to exploit this to find the extensions $X_{i}$ with uniformly bounded defect.